### v1

# Exact CHY Integrand Construction Using Combinatorial Neural Networks and Discrete Optimization

Abstract
We present a novel approach to the inverse CHY problem—constructing integrands from desired pole structures—by establishing a correspondence between CHY combinatorics and Combinatorial Neural Networks (CNNs). Our key innovation is the generalized pole degree $K(s_{A})$, which is additive under integrand multiplication and satisfies hierarchical recursion relations. These recursions naturally map to CNN message-passing architectures, enabling discrete automatic differentiation that preserves exact integer constraints throughout the computation. Unlike conventional machine learning approaches that yield numerical approximations, our method leverages the intrinsic combinatorial structure of the CHY formalism to guarantee exact algebraic solutions. This work demonstrates how specialized neural architectures can solve fundamental physics problems requiring strict mathematical constraints, providing a modest step toward bridging combinatorics, machine learning, and scattering amplitude theory.
Keywords:
Scattering Amplitudes, CHY Formalism, Combinatorial Neural Networks, Discrete Automatic Differentiation, Integer Linear Programming, Inverse Problem
1 Introduction
The Cachazo-He-Yuan (CHY) formalism Cachazo:2013gna ; Cachazo:2013iea ; Cachazo:2013hca provides a compact and universal representation of tree-level scattering amplitudes in a wide range of quantum field theories, including Yang-Mills, gravity, and scalar theories. In this framework, amplitudes are written as integrals over the moduli space of $n$ marked points on the Riemann sphere, localized by the scattering equations—a set of algebraic constraints that reflect momentum conservation and on-shell conditions.
While much progress has been made on evaluating CHY integrals for a given integrand—what we refer to as the “forward CHY problem”—a less explored and equally fundamental question is the inverse CHY problem: Given a desired set of physical singularities (i.e., pole structures), how does one construct a CHY integrand that produces them? This inverse direction is combinatorially complex, especially for high-multiplicity processes and integrands with nontrivial numerator structures.
In this paper, we propose a systematic approach to address the inverse CHY problem by establishing an unexpected connection between CHY integrands and combinatorial neural networks. We introduce a generalized pole degree $K(s_{A})$ that satisfies elegant recursion relations, transforming the inverse problem into integer linear programming. Most remarkably, these recursion relations naturally correspond to the architecture of combinatorial neural networks, enabling us to develop discrete automatic differentiation algorithms that guarantee exact solutions.
1.1 Background: The CHY Formalism and AI in Theoretical Physics
The CHY formalism expresses an $n$-particle scattering amplitude $\mathcal{A}_{n}$ as an integral over the positions of $n$ punctures $\{z_{i}\}$ on the Riemann sphere:
$\mathcal{A}_{n}=\int\frac{\prod_{i=1}^{n}dz_{i}}{\text{vol}(\mathrm{SL}(2,\mathbb{C}))}\;\Omega(\mathcal{E})\;\mathcal{I}(z).$ (1)
The measure $\Omega(\mathcal{E})$ enforces the scattering equations
$\mathcal{E}_{a}\equiv\sum_{b\neq a}\frac{s_{ab}}{z_{a}-z_{b}}=0,$ (2)
which relate external kinematics to the geometry of the punctured sphere. Different quantum field theories are encoded by different CHY integrands $\mathcal{I}(z)$, which specify the dynamical content of the amplitude. The formalism was originally introduced for tree-level amplitudes in a variety of massless theories Cachazo:2013hca ; Cachazo:2014xea , and later extended to polynomial forms Dolan:2014dia .
The CHY formalism’s remarkable universality has established it as a cornerstone of modern amplitude methods, providing a unifying framework that elegantly connects diverse quantum field theories. Since its inception, the formalism has revealed profound connections between seemingly disparate theories—from bi-adjoint scalars to Yang-Mills and gravity—through the simple modification of integrand functions $\mathcal{I}(z)$. This universality continues to yield new insights, as demonstrated by recent discoveries of novel factorization properties in Yang-Mills amplitudes He:2025a ; Zhang:2025 , which further illuminate the hidden structures within the CHY framework.
The formalism’s reach extends beyond tree-level physics, with significant progress in loop-level generalizations. Recent developments have successfully extended CHY methods to one-loop amplitudes in gauge theories Abreu:2025 . These advances build upon earlier breakthroughs, including the forward-limit construction for loop integrands He:2015yua ; Cachazo:2015nwa and the development of ambitwistor string theory connections Mason:2013sva ; Geyer:2015bja .
Within this rich landscape, computational tools have evolved to match the formalism’s sophistication. The “pick pole” algorithm, designed to isolate contributions containing specific physical singularities $1/s_{A}$, exemplifies the deep interplay between mathematical structure and practical computation Feng:2016nsv ; Feng:2019winn . For bi-adjoint scalar theory, this algorithm admits an elegant implementation based on contiguity properties of particle orderings. The extension to more general CHY integrands, particularly those associated with labelled tree graphs Gao:2017dek and higher-order poles Cardona:2016gon , reveals increasingly intricate mathematical structures that continue to inspire new approaches.
The computational challenges inherent in CHY calculations—from solving scattering equations to managing combinatorial complexity—have stimulated cross-disciplinary innovations. Recent work has explored connections to tropical geometry Cachazo:2019JHEP06 , positive geometries Arkani-Hamed:2017tmz , and intersection theory Mizera:2019 , demonstrating how the CHY formalism serves as a bridge between theoretical physics and modern mathematics. It is within this vibrant and expanding research landscape that we contribute our perspective, leveraging emerging techniques from discrete optimization and topological data analysis to address longstanding computational challenges while revealing new mathematical structures within the CHY framework.
Concurrent with these developments, theoretical physics has witnessed a paradigm shift with the integration of artificial intelligence and machine learning techniques. In string theory, machine learning has been employed to navigate the vast landscape of Calabi-Yau manifolds he2017 ; he2021 ; ashmore2020 . For scattering amplitudes, neural networks have shown promise in evaluating loop integrals Winterhalder:2022a ; Badger:2023 , optimizing phase-space integration Bothmann:2020 , and learning amplitude symmetries Alnuqaydan:2022 ; Dersy:2023 . The transformer architecture has proven remarkably effective for symbolic computations in quantum field theory cai2024 ; Alnuqaydan:2023symba . This progress reflects the emergence of AI-driven discovery as a new paradigm in science hey2009 ; xu2021 , where the goal extends beyond predictive models to developing explainable AI that translates patterns into physical insights carleo2019 ; Lucie-Smith:2020ris .
However, most machine learning applications in physics rely on continuous optimization yielding approximate results. For the CHY formalism, where exact algebraic structures and integer constraints are paramount, conventional approaches are unsuitable. This motivates our development of a fundamentally different framework based on discrete optimization and exact algorithms.
1.2 From Graphs to Combinatorial Neural Networks
Our key insight is that the pole structure of CHY integrands naturally exhibits a hierarchical organization that can be precisely captured by combinatorial neural networks—architectures originally developed for processing higher-order topological structures in data science Ebli:2020 ; Bodnar:2021 . This represents a conceptual leap: rather than applying standard machine learning tools to approximate solutions, we identify an exact correspondence between the mathematical structure of CHY integrands and these specialized neural architectures.
We begin by extending the conventional pole index $\chi(A)=L[A]-2(|A|-1)$ to introduce the generalized pole degree $K(s_{A})$, which measures the net number of internal edges within particle subsets. This quantity exhibits remarkable combinatorial properties: it behaves additively under integrand multiplication, transforming the complex problem of integrand composition into simple integer arithmetic. More importantly, $K(s_{A})$ satisfies hierarchical recursion relations:
$K(A)=\frac{1}{(|A|-2)!}\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-1\end{subarray}}K(B),$ (3)
which encode the complete pole structure in terms of two-particle data.
Building on these recursive properties, we formulate the inverse CHY problem as a systematic bootstrap procedure. Given a desired pole structure, we show that the construction of appropriate CHY integrands reduces to solving a mixed integer linear programming problem. The fundamental insight is that the hierarchical recursion relations create a system of linear constraints on the two-particle pole degrees $\{K(s_{ij})\}$, enabling systematic exploration of the space of possible theories within the CHY framework.
Interestingly, we discover that the hierarchical recursion structure of $K(s_{A})$ corresponds precisely to the architecture of combinatorial neural networks. This connection reveals that the inverse CHY problem can be reformulated as a discrete backpropagation process operating on simplicial complexes. Unlike continuous automatic differentiation used in standard machine learning frameworks, our approach requires integer discrete automatic differentiation that respects the combinatorial constraints of the CHY formalism. The algorithm learns the discrete structure of the CHY integrand space, navigating it efficiently to find solutions that precisely satisfy all specified pole requirements.
Leveraging this structure, we develop a neural-inspired algorithm that propagates constraints through the simplicial complex and enables effective bootstrapping of integrands. The algorithm bears resemblance to message-passing procedures in combinatorial neural networks and supports modular construction of integrands with higher-order poles. To validate our method, we analyze six- and eight-point examples, showing that our approach leads to more compact decompositions than traditional symmetric methods Cardona:2016gon .
Our framework not only provides a new perspective on the CHY landscape but also establishes a paradigm for applying machine learning to fundamental physics problems. By identifying deep structural correspondences between physics and computational frameworks, we demonstrate that certain problems naturally map to specialized architectures. As the field progresses toward more complex calculations at higher multiplicities and loop orders, such interdisciplinary approaches that leverage structural insights from both physics and machine learning will likely become increasingly important.
2 Graph-Theoretic Structure of CHY Integrands
2.1 Integrand as 4-Regular Graph and Integration Rules
The CHY formalism allows for a direct translation from the algebraic form of an integrand to a graph-theoretic object. For an $n$-particle process, the punctures $z_{i}$ are treated as vertices of a graph. For each factor of $(z_{i}-z_{j})^{-k}$ in the denominator of the integrand, $k$ edges are drawn between vertices $i$ and $j$. For the BAS integrand, $\mathcal{I}_{\text{BAS}}=\text{PT}(\alpha)\text{PT}(\beta)$, each PT factor contributes edges between cyclically adjacent vertices in its respective ordering. Since each vertex $k$ is connected to two neighbors in the $\alpha$ ordering and two neighbors in the $\beta$ ordering, every vertex has exactly four edges connected to it. Such a graph, where every vertex has a degree of 4, is known as a 4-regular (or quartic) graph Baadsgaard:2015voa ; Mason:2013sva ; Adamo:2013tsa ; Bern:2008qj ; Baadsgaard:2015aha ; Cachazo:2014nsa .
Figure 1: The 4-regular graph of integrand $\mathcal{I}=\text{PT}(123456)\text{PT}(125436)$
This graphical representation leads to a powerful set of ”integration rules,” which provide a purely combinatorial algorithm for evaluating the CHY integral without explicitly solving the scattering equations Baadsgaard:2015voa ; Baadsgaard:2015aha ; Cachazo:2014nsa ; Cardona:2016gon . The final amplitude is a sum over terms, each being a product of inverse Mandelstam variables, $\prod 1/s_{A_{i}}$. The rules determine which sets of Mandelstam variables appear.
  1. 1.
    Identify Potential Poles: For any subset of external labels $A\subset\{1,\dots,n\}$, a ”pole index” $\chi(A)$ is calculated:
    $\chi(A):=L[A]-2(|A|-1)$ (4)
[…]
  2. 2.
    Find Compatible Sets: A single term in the final amplitude corresponds to a valid Feynman diagram with $n-3$ propagators. The algorithm requires finding sets of $n-3$ distinct subsets $\{A_{1},\dots,A_{n-3}\}$, each satisfying $\chi(A_{i})=0$, that are ”compatible.” Compatibility is a graph-theoretic condition ensuring that the corresponding propagators can coexist in a single cubic Feynman diagram Geyer:2015bja .
  3. 3.
    Sum Contributions: Each compatible set contributes a term $1/(s_{A_{1}}\cdots s_{A_{n-3}})$ to the amplitude. The full amplitude is the sum over all such contributions Adamo:2013tsa .
For example, for a 6-point BAS amplitude with integrand $\mathcal{I}=\text{PT}(123456)\text{PT}(125436)$, one would first draw the corresponding 12-edge, 6-vertex, 4-regular graph. Then, one would identify all subsets $A$ with $\chi(A)=0$ (e.g., $A=\{1,2,6\}$) and find all compatible triplets of such subsets. The sum of the inverse products of the corresponding Mandelstam variables gives the final amplitude, matching the result from a traditional Feynman diagram calculation Bern:2008qj ; Baadsgaard:2015voa ; Eberhardt:2019qcl .
2.2 Combinatorial Structure of Generalized CHY Graphs
The graph-theoretic methods outlined for the bi-adjoint scalar (BAS) theory can be extended into a more general and powerful framework. This allows for a unified combinatorial analysis of any theory described by a rational CHY integrand.
First, we generalize the graph representation itself. To accommodate any rational function of differences $(z_{i}-z_{j})$, we introduce two types of edges:
  • A factor of $(z_{i}-z_{j})^{-1}$ in the denominator corresponds to a solid edge.
  • A factor of $(z_{i}-z_{j})^{+1}$ in the numerator corresponds to a dashed edge.
This establishes a direct map between any rational integrand $I(z)$ and a corresponding graph with two edge types.
With this representation, we can define a local property for each vertex $v$, its Net Degree, which counts the net power of factors involving $z_{v}$ in the integrand:
$\text{deg}_{\text{net}}(v):=\text{deg}_{\text{solid}}(v)-\text{deg}_{\text{dashed}}(v)$ (5)
This concept of a local net count can be conceptually extended to a ”regional” property for any subset of vertices $A\subset\{1,\dots,n\}$. This gives rise to the Generalized Pole Degree, K(A), which measures the net number of internal edges within the region defined by $A$:
$K(A):=L_{\text{solid}}[A]-L_{\text{dashed}}[A]$ (6)
where $L_{\text{solid}}[A]$ and $L_{\text{dashed}}[A]$ are the number of solid and dashed edges, respectively, with both endpoints in the set $A$. This new degree directly relates to the original pole index $\chi(A)$ found in the literature via the simple formula:
$\chi(A)=K(A)-2(|A|-1)$ (7)
[…]
The true power of this framework is revealed when we consider the multiplication of integrands, as in the ”double copy” construction. The multiplication of two graphs, $G_{\alpha}$ and $G_{\beta}$, corresponding to integrands $I_{\alpha}(z)$ and $I_{\beta}(z)$, is defined as a new graph $G_{\gamma}=G_{\alpha}*G_{\beta}$ corresponding to the product integrand $I_{\alpha}(z)I_{\beta}(z)$. The edge set of the product graph is the union of the edge sets of the original graphs. The index $K(s_{A})$ behaves very simply under this operation.
Theorem: The $K(s_{A})$ is additive under graph multiplication. Proof: For a product graph $G_{\gamma}=G_{\alpha}*G_{\beta}$, the number of internal edges in a subset $A$ is the sum of internal edges from the constituent graphs. Thus,
$K_{\gamma}(A)$ $:=L_{\text{solid, }\gamma}[A]-L_{\text{dashed, }\gamma}[A]$
$=(L_{\text{solid, }\alpha}[A]+L_{\text{solid, }\beta}[A])-(L_{\text{dashed, }\alpha}[A]+L_{\text{dashed, }\beta}[A])$
$=(L_{\text{solid, }\alpha}[A]-L_{\text{dashed, }\alpha}[A])+(L_{\text{solid, }\beta}[A]-L_{\text{dashed, }\beta}[A])$
$=K_{\alpha}(A)+K_{\beta}(A).\quad\square$
This additivity provides a simple algebraic tool to compute the pole structure of complicated theories by summing the known indices of their simpler constituents.
Figure 2: The ”0-regular graphs” is closed under multiplication
  $I^{[1]}$ $I^{[2]}$ $I^{[3]}$     $I^{[1]}$ $I^{[2]}$ $I^{[3]}$
    $deg(s_{1})$ 0 0 0     $K(s_{35})$ 1 -1 0
    $deg(s_{2})$ 0 0 0     $K(s_{36})$ 0 0 0
    $deg(s_{3})$ 0 0 0     $K(s_{45})$ 0 1 1
    $deg(s_{4})$ 0 0 0     $K(s_{46})$ 0 0 0
    $deg(s_{5})$ 0 0 0     $K(s_{56})$1 0 -1
    $deg(s_{6})$ 0 0 0     $K(s_{123})$ 0 0 0
    $K(s_{12})$ 1 -1 0     $K(s_{124})$ 1 0 1
    $K(s_{13})$ 0 0 0     $K(s_{125})$ 1 -1 0
    $K(s_{14})$ 0 1 1     $K(s_{126})$ 0 -1 -1
    $K(s_{15})$ 0 0 0     $K(s_{134})$ 1 -1 0
    $K(s_{16})$1 0 -1     $K(s_{135})$ 0 0 0
    $K(s_{23})$1 1 0     $K(s_{136})$1 1 0
    $K(s_{24})$ 0 0 0     $K(s_{145})$1 2 1
    $K(s_{25})$ 0 0 0     $K(s_{146})$1 1 0
    $K(s_{26})$ 0 0 0     $K(s_{156})$ 0 -1 -1
    $K(s_{34})$ 1 -2 -1
Figure 3: The $K_{I^{[3]}}(s_{A})=K_{I^{[1]}}(s_{A})+K_{I^{[2]}}(s_{A})$ for all poles
Figure 4: The ”0-regular group” acts as a set of transformations on other graphs
  $I^{[4]}$ $I^{[5]}$ $I^{[6]}$     $I^{[4]}$ $I^{[5]}$ $I^{[6]}$
    $deg(s_{1})$ 4 0 4     $K(s_{35})$ 0 -1 -1
    $deg(s_{2})$ 4 0 4     $K(s_{36})$ 1 0 1
    $deg(s_{3})$ 4 0 4     $K(s_{45})$ 2 1 3
    $deg(s_{4})$ 4 0 4     $K(s_{46})$1 0 -1
    $deg(s_{5})$ 4 0 4     $K(s_{56})$ 2 0 2
    $deg(s_{6})$ 4 0 4     $K(s_{123})$ 3 0 3
    $K(s_{12})$ 2 -1 1     $K(s_{124})$ 3 0 3
    $K(s_{13})$ 0 0 0     $K(s_{125})$ 2 -1 1
    $K(s_{14})$ 1 1 2     $K(s_{126})$ 4 -1 3
    $K(s_{15})$ 0 0 0     $K(s_{134})$ 3 -1 2
    $K(s_{16})$ 1 0 1     $K(s_{135})$ 0 0 0
    $K(s_{23})$ 1 1 2     $K(s_{136})$ 2 1 3
    $K(s_{24})$ 0 0 0     $K(s_{145})$ 3 2 5
    $K(s_{25})$ 0 0 0     $K(s_{146})$ 1 1 2
    $K(s_{26})$ 1 0 1     $K(s_{156})$ 3 -1 2
    $K(s_{34})$ 2 -2 0
Figure 5: The $K_{I^{[6]}}(s_{A})=K_{I^{[4]}}(s_{A})+K_{I^{[5]}}(s_{A})$ for all poles
[…]
2.3 Recursion Relation for Generalized Pole Degree $K(s_{A})$
The generalized pole degree $K(s_{A})$ introduced in the previous section exhibits a remarkable recursive structure that underlies the combinatorial organization of CHY integrands. This recursion provides a systematic way to compute pole degrees for arbitrary subsets of external particles from simpler building blocks.
For any non-empty subset $A\subseteq\{1,\ldots,n\}$ with cardinality $|A|\geq 3$, the generalized pole degree satisfies the fundamental recursion
$K(A)=\frac{1}{(|A|-2)!}\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-1\end{subarray}}K(B).$ (8)
This relation expresses the pole degree of any subset as a normalized sum over all its codimension-1 faces, revealing a hierarchical structure in the pole content of CHY integrands. The proof is given in Appendix A.
Illustrative examples.
  • Three points. For $s_{123}$ the recursion reads
    $K(s_{123})={K(s_{12})+K(s_{13})+K(s_{23})}.$
  • Four points. For $s_{1234}$ the recursion reads
    $K(s_{1234})=\frac{K(s_{123})+K(s_{124})+K(s_{134})+K(s_{234})}{2}.$
Figure 6: The recursion relation of $K(s_{A})$ on $I_{8}=\frac{z_{68}^{2}}{z_{13}^{2}z_{16}z_{17}z_{24}^{2}z_{28}^{2}z_{35}z_{36}z_{46}^{2}z_{56}z_{58}^{2}z_{67}z_{78}^{2}}$
The fundamental recursion (8) is a special case of a more general relation that connects pole degrees at different codimensions. For any integer $1\leq d\leq|A|-1$, we have
$K(A)=\frac{1}{\binom{|A|-2}{d}}\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-d\end{subarray}}K(B).$ (9)
The case $d=1$ recovers the original recursion (8). The proof is given in Appendix A.
The existence of multiple recursion relations at different codimensions provides valuable consistency checks for numerical implementations. For instance, when computing $K(12345)$, one can use: (i) $d=1$: sum over 5 four-point values with weight 1/61/61 / 6; (ii) $d=2$: sum over 10 three-point values with weight 1/31/31 / 3; (iii) $d=3$: sum over 10 two-point values with weight 1. Each route must yield the same result, providing a powerful diagnostic for computational accuracy.
[…]
2.4 Bootstrap Equations for Integrands
The relationship between the generalized pole degree $K(s_{A})$ and the pole index $\chi(A)$ established in equation (7) provides a powerful framework for solving the inverse CHY problem. This inverse problem asks: given a desired set of poles in a scattering amplitude, can we construct a CHY integrand that produces exactly those poles?
From equation (7), we know that the pole index is determined by (7). The fundamental principle of CHY integration rules tells us that only subsets $A$ with $\chi(A)=0$ contribute simple poles $1/s_{A}$ to the final amplitude Baadsgaard:2015voa ; Baadsgaard:2015aha . This immediately gives us a classification scheme for any subset $A\subset\{1,\dots,n\}$:
  • Physical poles: $\chi(A)=0$rightarrow simple poles $1/s_{A}$ appear in the amplitude
  • Absent poles: $\chi(A)<0$rightarrow no poles appear for these channels
  • Higher-order poles: $\chi(A)>0$rightarrow would correspond to multiple poles, but these typically do not appear in tree-level amplitudes for generic theories
This classification leads to a key insight: for any desired amplitude with a specific pole structure, the pattern of which poles appear and which are absent constrains the possible values of $K(s_{A})$. Consider a target amplitude that has simple poles for a set $\mathcal{P}=\{A_{1},A_{2},\ldots,A_{k}\}$ and no poles for the complementary set $\overline{\mathcal{P}}$. The bootstrap equations are:
$\text{For }A\in\mathcal{P}:\quad$ $K(A)=2(|A|-1)$ (10)
$\text{For }A\in\overline{\mathcal{P}}:\quad$ $K(A)<2(|A|-1)$ (11)
The equality constraints ensure that the desired poles appear with the correct order, while the inequality constraints ensure that unwanted poles are absent.
However, the values $K(A)$ cannot be chosen arbitrarily—they must satisfy the recursion relations derived in Section 2.3. This creates a highly constrained system where the fundamental unknowns are the two-particle pole degrees $K(s_{ij})$ for all pairs $\{i,j\}$. Using the recursion relation (8), any higher-order pole degree can be expressed in terms of lower-order ones. For example:
$K(s_{ijk})$ $=K(s_{ij})+K(s_{ik})+K(s_{jk})$ (12)
$K(s_{ijkl})$ $=\frac{K(s_{ijk})+K(s_{ijl})+K(s_{ikl})+K(s_{jkl})}{2}$ (13)
Continuing this process, we can express all pole degrees $K(A)$ with $|A|\geq 3$ as linear combinations of the two-particle pole degrees $K(s_{ij})$. This reduces our bootstrap system to a set of linear constraints on the $\binom{n}{2}$ variables $\{K(s_{ij})\}_{1\leq i<j\leq n}$.
A crucial observation is that for rational CHY integrands constructed from products of Parke-Taylor factors and their generalizations, all pole degrees $K(s_{A})$ are integers. This follows from the fact that each factor $(z_{i}-z_{j})^{\pm 1}$ contributes ±1$\pm$ 1± 1 to the edge count. Therefore, the bootstrap equations become a mixed system of linear Diophantine equations and inequalities:
$\mathbf{M}_{\text{eq}}\vec{K}_{2}$ $=\vec{b}_{\text{eq}}$ (14)
$\mathbf{M}_{\text{ineq}}\vec{K}_{2}$ $<\vec{b}_{\text{ineq}}$ (15)
where $\vec{K}_{2}=(K(s_{12}),K(s_{13}),\ldots,K(s_{n-1,n}))^{T}$ is the vector of two-particle pole degrees, $\mathbf{M}_{\text{eq}}$ and $\mathbf{M}_{\text{ineq}}$ encode the recursion relations for equality and inequality constraints respectively, while $\vec{b}_{\text{eq}}$ and $\vec{b}_{\text{ineq}}$ encode the corresponding target pole structures.
The bootstrap approach yields a constructive algorithm for the inverse CHY problem. The basic idea is to transform the desired pole structure into a system of linear constraints on the fundamental variables $K(s_{ij})$, solve this constrained system, and then use the additivity property under graph multiplication to construct the corresponding integrand. The algorithm proceeds in the following steps:
Algorithm 1 CHY Integrand Bootstrap Construction
0: Target pole structure $\mathcal{P}$ (set of desired simple poles)
0: CHY integrand $I(z)$ producing amplitude with pole structure $\mathcal{P}$
1:Input: Specify desired pole structure $\mathcal{P}\subset\mathcal{S}_{n}$ where $\mathcal{S}_{n}$ is the set of all possible subsets
2:Constraint Setup: For each $A\in\mathcal{P}$, impose $K(A)=2(|A|-1)$
3:Constraint Setup: For each $A\in\overline{\mathcal{P}}$, impose $K(A)<2(|A|-1)$
4:Recursion Matrix: Construct matrices $\mathbf{M}_{\text{eq}},\mathbf{M}_{\text{ineq}}$ encoding relations between $K(A)$ and $K(s_{ij})$
5:Mixed Integer System: Formulate $\mathbf{M}_{\text{eq}}\vec{K}_{2}=\vec{b}_{\text{eq}}$ and $\mathbf{M}_{\text{ineq}}\vec{K}_{2}<\vec{b}_{\text{ineq}}$
6:Integer Solution: Solve for integer vector $\vec{K}_{2}$ satisfying the mixed system
7:if no integer solution exists then
8:  return  ”No CHY integrand exists for this pole structure”
9:end if
10:Decomposition: Use additivity $K_{\alpha\star\beta}=K_{\alpha}+K_{\beta}$ to factor solution
11:Construction: Build integrand $I(z)=\prod_{i}I_{i}(z)$ from elementary factors
12:return  CHY integrand $I(z)$
For illustration, consider a six-particle amplitude where we want simple poles for $s_{12}$, $s_{34}$, and $s_{56}$, but no other two-particle poles. The bootstrap equations become $K(s_{12})=K(s_{34})=K(s_{56})=0$ and $K(s_{ij})<0$ for all other pairs $\{i,j\}$. Combined with the recursion relations for higher-order poles, this system constrains the possible CHY integrands. One solution corresponds to the integrand $I=\frac{1}{z_{12}^{2}z_{14}z_{15}z_{23}z_{26}z_{34}^{2}z_{36}z_{45}z_{56}^{2}}$, which indeed produces poles only in the appropriate channels.
This bootstrap framework thus provides a systematic method to reverse-engineer CHY integrands from desired amplitude structures, opening new possibilities for amplitude construction and the exploration of novel theories within the CHY formalism. The method is particularly powerful because it leverages the additive structure of the generalized pole degree $K(s_{A})$, allowing complex integrands to be built systematically from simpler components.
2.5 Examples of Bootstrap Construction
To illustrate the power of the bootstrap framework developed in Section 2.4, we present concrete examples demonstrating how to reverse-engineer CHY integrands from desired pole structures. These examples showcase both the computational implementation of the bootstrap algorithm and the elegant interplay between the algebraic constraints and the resulting graph structures.
Consider a six-particle amplitude where we desire simple poles for the channels $s_{12}$, $s_{23}$, $s_{34}$, $s_{56}$, $s_{123}$, $s_{156}$, $s_{234}$, and $s_{456}$. This represents a highly constrained scenario with eight simple poles, approaching the maximum possible for a six-particle amplitude.
Following the bootstrap algorithm, we first translate these requirements into constraints on the K-values. The equality constraints (10) yield:
$K(s_{12})=K(s_{23})=K(s_{34})=K(s_{56})$ =2=2= 2 (16)
$K(s_{123})=K(s_{156})=K(s_{234})=K(s_{456})$ =4=4= 4 (17)
For all other channels, the inequality constraints (11) require $K(s_{A})<2(|A|-1)$. Additionally, the single-particle K-values must satisfy the regularity condition:
$K_{i}=\sum_{j\neq i}K(s_{ij})=4\quad\text{for all }i\in\{1,\ldots,6\}$ (18)
This yields a mixed integer linear programming problem with 15 fundamental variables $K(s_{ij})$, subject to 14 equality constraints (8 from desired poles, 6 from single-particle conditions) and 27 inequality constraints from unwanted channels. The constraint matrix $\mathbf{M}_{\text{eq}}$ encodes how each higher-order K-value depends on the two-particle degrees through the recursion relations. For instance, the constraint for $K(s_{123})=4$ translates to:
$K(s_{12})+K(s_{13})+K(s_{23})=4$ (19)
The complete system of inequalities includes 27 inequality constraints (from unwanted two-particle and three-particle channels):
$K_{1,3}<2,K_{1,4}<2,K_{1,5}<2,K_{1,6}<2,K_{2,4}<2,K_{2,5}<2,K_{2,6}<2,$
$K_{3,5}<2,K_{3,6}<2,K_{4,5}<2,K_{4,6}<2,K_{1,2}+K_{1,4}+K_{2,4}<4,$
$K_{1,2}+K_{1,5}+K_{2,5}<4,K_{1,2}+K_{1,6}+K_{2,6}<4,K_{1,3}+K_{1,4}+K_{3,4}<4,$
$K_{1,3}+K_{1,5}+K_{3,5}<4,K_{1,3}+K_{1,6}+K_{3,6}<4,K_{1,4}+K_{1,5}+K_{4,5}<4,$
$K_{1,4}+K_{1,6}+K_{4,6}<4,K_{2,3}+K_{2,5}+K_{3,5}<4,K_{2,3}+K_{2,6}+K_{3,6}<4,$
$K_{2,4}+K_{2,5}+K_{4,5}<4,K_{2,4}+K_{2,6}+K_{4,6}<4,K_{2,5}+K_{2,6}+K_{5,6}<4,$
$K_{3,4}+K_{3,5}+K_{4,5}<4,K_{3,4}+K_{3,6}+K_{4,6}<4,K_{3,5}+K_{3,6}+K_{5,6}<4,$
and the complete system of 14 equality constraints (8 from desired poles, 6 from single-particle conditions):
$K_{1,2}=2,K_{2,3}=2,K_{3,4}=2,K_{5,6}=2,K_{1,2}+K_{1,3}+K_{2,3}=4,$
$K_{1,5}+K_{1,6}+K_{5,6}=4,K_{4,5}+K_{4,6}+K_{5,6}=4,K_{2,3}+K_{2,4}+K_{3,4}=4,$
$K_{1}=K_{1,2}+K_{1,3}+K_{1,4}+K_{1,5}+K_{1,6}=4,$
$K_{2}=K_{1,2}+K_{2,3}+K_{2,4}+K_{2,5}+K_{2,6}=4,$
$K_{3}=K_{1,3}+K_{2,3}+K_{3,4}+K_{3,5}+K_{3,6}=4,$
$K_{4}=K_{1,4}+K_{2,4}+K_{3,4}+K_{4,5}+K_{4,6}=4,$
$K_{5}=K_{1,5}+K_{2,5}+K_{3,5}+K_{4,5}+K_{5,6}=4,$
$K_{6}=K_{1,6}+K_{2,6}+K_{3,6}+K_{4,6}+K_{5,6}=4.$
Solving this system yields the unique integer solution:
$K_{1,2}$ 2$\to$ 2,→ 2 , $K_{1,3}$ 0$\to$ 0,→ 0 , $K_{1,4}$ 0$\to$ 0,→ 0 , (20)
$K_{2,3}$ 2$\to$ 2,→ 2 , $K_{2,4}$ 0$\to$ 0,→ 0 , $K_{2,5}$ 0$\to$ 0,→ 0 ,
$K_{2,6}$ 0$\to$ 0,→ 0 , $K_{3,4}$ 2$\to$ 2,→ 2 , $K_{3,5}$ 0$\to$ 0,→ 0 ,
$K_{3,6}$ 0$\to$ 0,→ 0 , $K_{4,5}$ 1$\to$ 1,→ 1 , $K_{4,6}$ 1$\to$ 1,→ 1 ,
$K_{5,6}$ 2$\to$ 2,→ 2 , $K_{1,5}$ 1$\to$ 1,→ 1 , $K_{1,6}$ 1$\to$ 1→ 1
The resulting CHY integrand:
$I_{6}=\frac{1}{z_{12}^{2}z_{15}z_{16}z_{23}^{2}z_{34}^{2}z_{45}z_{46}z_{56}^{2}}$ (21)
demonstrates that the bootstrap construction successfully produces an integrand with precisely the desired pole structure. The resulting integrand $I_{6}$ generates an amplitude whose only simple poles are $s_{12}$, $s_{23}$, $s_{34}$, $s_{56}$, $s_{123}$, $s_{156}$, $s_{234}$, and $s_{456}$, exactly as specified in our requirements. The K-values directly translate to the powers in the integrand: each $K(s_{ij})=2$ yields a factor $z_{ij}^{-2}$ in the denominator, while $K(s_{ij})=1$ gives $z_{ij}^{-1}$, and $K(s_{ij})=0$ means the corresponding factor is absent.
For this particular example, the mixed integer linear programming problem yields a unique solution, making it an ideal illustration of the bootstrap method. The constraint system can be solved using standard integer programming packages available in Mathematica (through functions like LinearProgramming with integer constraints) or Python (using libraries such as cvxpy, PuLP, or scipy.optimize.milp). These computational tools efficiently handle the combinatorial nature of finding integer solutions to our linear constraint system. More complex examples with non-unique solutions and detailed computational methods for solving the bootstrap equations are presented in Appendix B.
The graph-theoretic framework developed in this section has revealed the deep combinatorial structure underlying CHY integrands. Starting from the observation that integrands can be represented as generalized graphs with solid and dashed edges, we discovered three fundamental principles:
Additive Structure: The generalized pole degree $K(s_{A})$ is additive under integrand multiplication, leading to a group structure on 0-regular graphs that acts transitively on integrand spaces. This algebraic property transforms the complex problem of integrand multiplication into simple integer addition.
Recursive Organization: The pole degrees satisfy elegant recursion relations that encode the full pole structure in the two-particle data.
Bootstrap Construction: The inverse problem—constructing integrands from desired poles—reduces to integer linear programming, providing a systematic approach to exploring new theories within the CHY framework and guaranteeing the existence and uniqueness of solutions for compatible pole structures.
These algebraic and combinatorial insights suggest a natural extension to higher-dimensional structures. The recursion relations, which connect pole degrees across different particle subsets, hint at an underlying topological organization where the pole hierarchy can be viewed as a simplicial complex. This perspective may offer additional mathematical tools for understanding the CHY formalism.
The discrete optimization problems arising from the bootstrap construction also suggest potential connections to computational methods from combinatorial optimization and machine learning. In particular, algorithms designed for discrete structures on graphs and simplicial complexes could potentially be adapted to address computational challenges in amplitude calculations.
In the next section, we explore this topological perspective more systematically. By viewing the pole structure as a simplicial complex rather than just a graph, we uncover additional mathematical structure that provides both theoretical insights and practical algorithms for the pole selection problem. This approach connects the CHY formalism to recent developments in topological data analysis and suggests new computational strategies for working with multi-particle amplitudes.
3 From Poles to Simplices: A Topological Perspective
3.1 Simplicial Structure in CHY Pole Hierarchies
The pole structure of CHY integrands exhibits a natural hierarchical organization that goes beyond the graph-theoretic framework presented in Section 2. This hierarchy is most clearly revealed when we consider the partial ordering of particle subsets by inclusion, which forms a Hasse diagram structure that is intimately connected to simplicial complexes. Consider the collection of all subsets $A\subseteq{1,\ldots,n}$ ordered by inclusion. For each subset $A$, we have the pole index $\chi(A)=K(A)-2(|A|-1)$ determining whether $s_{A}$ appears as a pole in the amplitude. The recursion relations (8) and (9) show that the pole degree $K(A)$ of any subset is completely determined by the pole degrees of its proper subsets. This creates a layered structure where:
[…]
$K(\sigma)=\frac{1}{(\dim(\sigma)-1)!}\sum_{\tau\prec_{1}\sigma}K(\tau)$ (22)
where $\tau\prec_{1}\sigma$ denotes that $\tau$ is a codimension-1 face of $\sigma$. This simplicial perspective transforms the CHY pole structure into a topological data type. The physical poles ($\chi(A)=0$) form a subcomplex, while the absent poles ($\chi(A)<0$) and the forbidden higher-order poles ($\chi(A)>0$) partition the remaining simplices. The bootstrap equations (10)-(11) become constraints on which simplices can appear in the ”physical subcomplex.” The passage from graphs to simplicial complexes also enables us to apply tools from combinatorial Hodge theory. Define the space of $k$-chains $C_{k}$ as formal linear combinations of $k$-simplices, with the boundary operator $\partial_{k}:C_{k}\to C_{k-1}$ given by the standard alternating sum over faces. The pole degree function $K$ extends to a cochain $K\in C^{*}$, and the recursion relations imply a cohomological structure:
$\delta K=\Omega$ (23)
where $\delta$ is the coboundary operator and $\Omega$ encodes the deviations from the naive dimensional counting. This reveals that the pole structure has both a ”local” contribution (from the coboundary) and a ”harmonic” contribution (from the kernel of the Laplacian $\Delta=\delta\partial+\partial\delta$). Furthermore, the multiplicative structure of CHY integrands under the double-copy construction translates into operations on simplicial complexes. The additivity of $K(A)$ under graph multiplication (Theorem in Section 2.3) extends to a functorial relationship between simplicial complexes, where the product of integrands corresponds to a fiber product of the associated complexes with weights adding fiberwise. This topological perspective provides powerful new tools for analyzing CHY integrands:
Persistent homology can track how the pole structure changes under deformations Spectral sequences can organize the computation of amplitudes Sheaf cohomology can capture global constraints on consistent pole configurations
The simplicial structure naturally suggests the use of Combinatorial Neural Networks (CNNs), a class of neural architectures designed to operate on higher-order topological structures beyond graphs ebli2020 ; Bodnar:2021 . Unlike traditional Graph Neural Networks (GNNs) that are limited by the Weisfeiler-Lehman hierarchy xu2019 ; morris2019 , CNNs can process information across all dimensions of the simplicial complex simultaneously. This is particularly relevant for CHY integrands, where the pole structure involves intricate relationships among multi-particle channels that cannot be captured by pairwise interactions alone.
Recent developments in topological deep learning Zia:2023tdl ; Papillon:2023tddl and geometric machine learning on simplicial complexes Giusti:2022san ; Maggs:2023nnk provide the theoretical foundation for these approaches. The ability of CNNs to perform message passing along both boundary and coboundary relations, and to aggregate information hierarchically across different dimensions, makes them a natural computational framework for the recursive pole structures we have identified. The CHY formalism, with its rich combinatorial and topological structure, presents an ideal testing ground for these emerging methods at the intersection of algebraic topology and deep learning.
In the following sections, we will develop this correspondence in detail, showing how the combinatorial and topological structures work together with modern machine learning techniques to provide both theoretical insights and practical computational advantages for the CHY formalism.
3.2 Combinatorial Neural Networks and Discrete Automatic Differentiation
The topological perspective developed in Section 3 naturally leads us to consider computational frameworks that can efficiently handle the hierarchical pole structures of CHY integrands. While traditional Graph Neural Networks (GNNs) have revolutionized many areas of physics Bronstein:2021gdl , they are fundamentally limited to pairwise interactions and cannot capture the higher‑order relationships inherent in simplicial complexes Bodnar:2021mpsn ; Ebli:2020snn . This is where Combinatorial Neural Networks (CNNs) and automatic differentiation techniques become essential tools for the CHY formalism.
A crucial observation is that the pole adjustment algorithm operates on discrete structures—the K-values are integers, and the operations involve discrete increments and decrements. This is fundamentally different from the continuous automatic differentiation implemented in popular frameworks like TensorFlow, PyTorch, or JAX paszke2017automatic ; abadi2016tensorflow ; bradbury2018jax . Current frameworks are designed for continuous optimization and lack native support for:
  • Integer-valued gradients and discrete optimization
  • Topological constraints on simplicial complexes
  • Exact satisfaction of linear Diophantine equations
The discrete nature of our problem requires a specialized automatic differentiation framework that can handle:
$\frac{\partial K(s_{A})}{\partial K(s_{ij})}\in\mathbb{Z}$ (24)
where the ”derivatives” are actually sensitivity coefficients derived from the recursion relations (8).
For a $k$-simplex $\sigma$ in our pole hierarchy, the CNN update rule takes the form Bodnar:2021 :
$h_{\sigma}^{(t+1)}=\phi\left(h_{\sigma}^{(t)},\bigoplus_{\tau\in\mathcal{B}(\sigma)}m_{\tau\to\sigma}^{(t)},\bigoplus_{\tau\in\mathcal{C}(\sigma)}m_{\tau\to\sigma}^{(t)}\right)$ (25)
where $\mathcal{B}(\sigma)$ denotes the boundary (codimension-1 faces) and $\mathcal{C}(\sigma)$ denotes the coboundary (simplices containing $\sigma$). This bidirectional information flow perfectly mirrors the upward and downward propagation we observe in the pole adjustment algorithm.
Algorithm 2 PropagateUpdate: Discrete Automatic Differentiation
0: Graph $G$, modified pole $s_{A}$, change $\Delta K$, fixed poles $\mathcal{F}$
0: Updated K-values satisfying all constraints
1:Queue $Q\leftarrow\{s_{A}\}$, Visited $V\leftarrow\emptyset$
2:while $Q\neq\emptyset$ do
[…]
4:  if $s_{B}\in V$ then
5:   continue
6:  end if
7:  $V\leftarrow V\cup\{s_{B}\}$
8:  {Upward propagation to supersets}
9:  for each superset $s_{C}\supset s_{B}$ with $|s_{C}|=|s_{B}|+1$ do
10:   if $s_{C}\notin\mathcal{F}$ then
11:    $\Delta K_{C}\leftarrow\Delta K/(|s_{B}|-1)!$
12:    $K^{*}(s_{C})\leftarrow K^{*}(s_{C})+\Delta K_{C}$
[…]
14:   else
15:    {Fixed point encountered, trigger backpropagation}
[…]
17:   end if
18:  end for
19:  {Backward propagation to subsets (if from higher layer)}
20:  if $|s_{B}|>2$ then
21:   for each subset $s_{D}\subset s_{B}$ with $|s_{D}|=|s_{B}|-1$ do
22:    if $s_{D}\notin\mathcal{F}$ then
23:     Select one $s_{D}$ and update (choice based on constraints)
24:     $K^{*}(s_{D})\leftarrow K^{*}(s_{D})+\Delta K$
[…]
26:    end if
27:   end for
28:  end if
29:end while
The key insight is that our ”gradients” are actually linear relationships derived from the recursion relations. For any modification $\Delta K(s_{ij})$ to a two-point pole, the change propagates through the hierarchy according to:
$\Delta K(s_{A})=\sum_{\begin{subarray}{c}i,j\in A\\ i<j\end{subarray}}c_{A,ij}\Delta K(s_{ij})$ (26)
where the coefficients $c_{A,ij}\in\mathbb{Q}$ are determined by the combinatorial structure of the simplex.
The training process for this specialized CNN architecture proceeds through two fundamental phases that work together to maintain the consistency of the pole hierarchy:
1. Upward Propagation Update: When we modify a pole degree $K(s_{A})$ by an integer amount $\Delta K$, this change must propagate upward through the hierarchy to all supersets to maintain the recursion relations. As illustrated in Figure 8, the algorithm systematically updates all affected higher-order poles according to the formula:
$K(s_{B})\leftarrow K(s_{B})+\frac{\Delta K}{(|A|-1)!}\quad\text{for all }B\supset A\text{ with }|B|=|A|+1$ (27)
2. Backward Propagation (Triggered by Fixed Nodes): When the upward propagation encounters a fixed pole—one whose value must remain constant due to physical constraints—the algorithm triggers a backward propagation phase. As shown in Figure 9, this phase redistributes the constraints by selecting appropriate subsets and adjusting their values to compensate for the fixed constraint while maintaining all recursion relations.
These two processes continue alternating until the Combinatorial Neural Network reaches a stable configuration where all constraints are satisfied. At convergence, we extract the bottom layer K-values (the two-particle pole degrees $\{K(s_{ij})\}$), which completely determine the CHY integrand through:
$I(z)=\prod_{1\leq i<j\leq n}(z_{i}-z_{j})^{-K(s_{ij})}$ (28)
[…]
Figure 8: Upward propagation in the discrete automatic differentiation algorithm. (a) Initial configuration showing the two-particle pole $s_{13}$ with $K(s_{13})=2$ being decremented by 1 (green node). The gray edges represent the dependency relationships from the recursion relation $K(s_{ABC})=K(s_{AB})+K(s_{AC})+K(s_{BC})$. (b) After applying $\Delta K=-1$ to $s_{13}$, the change propagates upward to all three-particle supersets containing particles 1 and 3. Each superset $\{s_{123},s_{134},s_{135},s_{136}\}$ receives an update of 1-1- 1 (shown as blue edges with -1 labels), maintaining the recursion relations. The updated values are shown in cyan nodes.
[…]
Figure 9: Backward propagation triggered by fixed constraints in the pole hierarchy. (a) Initial state with fixed poles $s_{12}$, $s_{123}$, $s_{2}$, and $s_{3}$ (red nodes) that must maintain their values throughout the algorithm. The pole $s_{23}$ (green) is to be decremented by 1. (b) Upward propagation attempts to update the three-particle poles, but encounters the fixed constraint at $s_{123}$ (and implicitly at single-particle nodes $s_{2}$ and $s_{3}$). (c) Backward propagation is triggered to maintain consistency: to compensate for the inability to modify fixed nodes, the algorithm selects $s_{13}$ (intersection of subsets from $s_{123}$ and supersets of $s_{3}$) and $s_{24}$ (superset of $s_{2}$), incrementing each by +1 (red edges). This redistribution ensures all recursion relations remain satisfied while respecting the fixed constraints.
This discrete automatic differentiation ensures:
  • All K-values remain integers (when changes are properly scaled)
  • Recursion relations are exactly satisfied at each step
  • Fixed poles maintain their values throughout propagation
The backpropagation step, triggered when encountering fixed points, implements a discrete version of the chain rule that respects the topological constraints of the simplicial complex. This is fundamentally different from continuous backpropagation and requires specialized algorithms not found in current deep learning frameworks griewank2008evaluating .
The examples in the following sections serve both as validation of the algorithm and as intuitive illustrations of how discrete automatic differentiation operates on the CHY pole hierarchy.
3.3 Neural-Inspired Pick-Pole Algorithm
The computational framework developed in Section 3.2, combining Combinatorial Neural Networks with discrete automatic differentiation, provides the essential infrastructure for implementing the pole selection algorithm. We now turn to the central computational challenge in the CHY formalism: isolating specific physical singularities from the complete amplitude structure.
The pick-pole algorithm addresses a fundamental problem in the practical application of the CHY formalism Feng:2016nsv ; Feng:2019winn . Given a CHY integrand $I(z)$ that produces an amplitude with multiple poles, we often need to isolate contributions containing a specific pole $1/s_{A}$ where $s_{A}=(\sum_{i\in A}k_{i})^{2}$. This capability is essential for analyzing the analytic structure of amplitudes, studying factorization limits when particles become collinear or soft, and constructing loop-level integrands from tree-level data via forward limits and unitarity cuts He:2015yua ; Cachazo:2015nwa .
For the bi-adjoint scalar theory, where the integrand takes the form $I[\pi|\rho]=\text{PT}(\pi)\times\text{PT}(\rho)$, the algorithm is elegant and provably correct Baadsgaard:2015voa ; Huang:2018pbu . The key insight is that a pole $s_{A}$ appears if and only if the set $A$ forms a contiguous block in both orderings $\pi$ and $\rho$. One can then multiply by an appropriate cross-ratio factor to eliminate unwanted contributions:
$\mathcal{P}_{bd}^{ac}=\frac{z_{ac}z_{bd}}{z_{ad}z_{cb}}$ (29)
where $a,b$ are the endpoints of the contiguous set $A$ and $c,d$ are their nearest neighbors in the complement $\bar{A}$ Feng:2020lgt .
For more general CHY integrands, particularly those associated with labelled tree graphs Gao:2017dek , the notion of “contiguity” loses meaning, and the algorithm becomes significantly more complex. As noted in Feng:2020lgt , the current approach is fundamentally heuristic, relying on a generate-and-test search through combinations of cross-ratio factors. This heuristic method, while powerful enough to construct the first universal one-loop CHY integrand for BAS theory Feng:2019winn , suffers from exponential computational complexity and lacks a rigorous proof of correctness.
Our approach reformulates this challenge within the graph-theoretic and topological framework developed in previous sections. The key insight is that pole selection corresponds to a systematic adjustment of K-values while maintaining all recursion relations and fixed-point constraints. Rather than searching through combinations of cross-ratio factors, we directly manipulate the underlying graph structure through controlled modifications of the generalized pole degrees.
Algorithm 3 Pick-Pole Algorithm for CHY Integrands
0: CHY integrand $I(z)$ with pole structure, target pole $s_{T}$ to retain
0: Modified integrand $I^{*}(z)$ containing only terms with pole $s_{T}$
1:Initialize: Construct pole hierarchy graph $G$ with K-values from $I(z)$
2:Identify: Mark poles to retain $\mathcal{P}_{\text{retain}}=\{s_{T}\}\cup\text{CompatiblePoles}(s_{T})$
3:Identify: Mark poles to remove $\mathcal{P}_{\text{remove}}=\text{AllPoles}(I)\setminus\mathcal{P}_{\text{retain}}$
4:for each pole $s_{A}\in\mathcal{P}_{\text{remove}}$ do
5:  $K^{*}(s_{A})\leftarrow K(s_{A})-1$Decrement K-value}
[…]
7:end for
8:return  Integrand $I^{*}(z)$ constructed from modified K-values
The algorithm’s correctness relies on several key properties established in previous sections. The additivity of K-values (Theorem in Section 2.2) ensures that modifications propagate correctly through the graph structure. The recursion relations (8)-(9) maintain consistency across different levels of the pole hierarchy. Most critically, the discrete automatic differentiation framework embodied in the PropagateUpdate subroutine (Algorithm 2) ensures that all constraints are satisfied throughout the modification process.
The connection to the topological perspective is particularly illuminating. Pole selection can be viewed as a projection operation on the simplicial complex, where we retain only those simplices (poles) compatible with our target. This geometric interpretation provides both intuition and a correctness criterion: the resulting integrand must correspond to a valid subcomplex of the original pole structure. The discrete nature of our approach—working with integer-valued K-degrees and exact constraint satisfaction—distinguishes it fundamentally from continuous optimization methods and ensures that the output is always a valid CHY integrand.
The following sections demonstrate the algorithm’s effectiveness through explicit examples, showing how it successfully isolates desired pole structures in both six-point and eight-point amplitudes. These examples serve not only as validation of the method but also as pedagogical illustrations of how the discrete automatic differentiation operates on the hierarchical pole structure of CHY integrands.
3.4 Examples of 6-Point CHY Integrands
We demonstrate the neural-inspired pick-pole algorithm through a concrete 6-point example. The algorithm operates on discrete structures where K-values are integers, and the pole hierarchy forms a two-layer network structure that enables systematic pole selection through discrete automatic differentiation.
For any subset $A\subset\{1,2,\cdots,n\}$, we define the K-value as:
$K(s_{A}):=L[A]$
Consider a six-point CHY integrand:
$I_{6}=\frac{1}{z_{12}^{2}z_{23}^{2}z_{34}^{2}z_{56}^{2}z_{16}z_{45}z_{46}z_{15}}$
which produces five cubic Feynman diagrams:
$\frac{1}{s_{12}s_{34}s_{56}}+\frac{1}{s_{123}s_{12}s_{56}}+\frac{1}{s_{123}s_{23}s_{56}}+\frac{1}{s_{156}s_{23}s_{56}}+\frac{1}{s_{156}s_{34}s_{56}}$
To pick pole $s_{12}$, we need to retain poles $s_{12}$, $s_{34}$, $s_{56}$ and $s_{123}$ while removing poles $s_{23}$ and $s_{156}$.
Figure 10: Initial pole hierarchy for 6-point CHY integrand : Two-layer network structure with two-particle poles in the first layer (bottom), single-particle and three-particle poles in the second layer (top). Gray edges represent dependency relationships from recursion relation $K(s_{abc})=K(s_{ab})+K(s_{ac})+K(s_{bc})$ and $deg(s_{a})=K(s_{ab})+K(s_{ac})+K(s_{ad})+K(s_{ae})+K(s_{af})$. Red nodes indicate fixed poles that must be preserved during the algorithm.
The recursion relations defining the pole hierarchy are:
$K(s_{abc})=K(s_{ab})+K(s_{ac})+K(s_{bc})$
$deg(s_{a})=K(s_{ab})+K(s_{ac})+K(s_{ad})+K(s_{ae})+K(s_{af})$
We fix the K-values of desired poles:
$K(s_{12})=2,\quad K(s_{34})=2,\quad K(s_{56})=2,\quad K(s_{123})=4$
Step 1: Remove pole $s_{23}$
We decrement the K-value of the unwanted pole $s_{23}$:
$K^{*}(s_{23})=K(s_{23})-1=1$
Figure 11: First upward propagation : The modification $\Delta K=-1$ at pole $s_{23}$ (green node) propagates upward through the network hierarchy. Blue arrows show the upward propagation paths to single-particle subsets and three-particle supersets. The algorithm encounters fixed constraints at $s_{123}$, $s_{2}$, and $s_{3}$ (red nodes), triggering the backward propagation mechanism.
The upward propagation encounters fixed points $s_{123}$, $s_{2}$, and $s_{3}$, triggering backward propagation.
Figure 12: First backward propagation : When upward propagation hits fixed constraints, the discrete automatic differentiation algorithm redistributes the changes through backward propagation (red arrows). The algorithm selects intersection $s_{13}$ and superset $s_{24}$, incrementing their K-values by +1 to maintain all recursion relations while respecting fixed constraints.
The algorithm compensates by selecting appropriate poles:
$K^{*}(s_{13})=K(s_{13})+1=1,\quad K^{*}(s_{24})=K(s_{24})+1=1$
Figure 13: Second upward propagation : The changes at $s_{13}$ and $s_{24}$ propagate upward (blue arrows), updating their three-particle supersets $s_{134}$, $s_{135}$, $s_{136}$ and $s_{124}$ by +1. The propagation again encounters fixed points $s_{1}$ and $s_{4}$, necessitating another round of backward propagation.
The upward propagation again encounters fixed points $s_{1}$ and $s_{4}$.
Figure 14: Second backward propagation : To maintain consistency with fixed single-particle constraints, the algorithm performs backward propagation from the second layer to the first layer (red arrows). The intersection $s_{14}$ of dependent supersets from fixed nodes $s_{1}$ and $s_{4}$ is decremented by -1, ensuring all recursion relations remain satisfied.
Figure 15: Final upward propagation for $s_{23}$ removal : The third upward propagation completes without encountering additional fixed points, indicating successful convergence of the discrete automatic differentiation algorithm. The removal of pole $s_{23}$ is now complete with all constraints satisfied.
The algorithm terminates when no fixed points are encountered, completing the removal of pole $s_{23}$.
Step 2: Remove pole $s_{156}$
We now repeat the procedure to remove pole $s_{156}$.
Figure 16: Updated pole hierarchy for $s_{156}$ removal : The network structure after removing $s_{23}$, showing updated K-values throughout the hierarchy. Fixed points (red) and the target pole $s_{156}$ for removal (green) are clearly marked. The algorithm now proceeds to remove the second unwanted pole while maintaining all established constraints.
We decrement the K-value of pole $s_{156}$:
$K^{*}(s_{156})=K(s_{156})-1=3$
Figure 17: Downward propagation : Unlike the previous case, the modification at three-particle pole $s_{156}$ requires downward propagation to the first layer (red arrows). The algorithm selects subset $s_{15}$ and decrements its K-value by -1, demonstrating the bidirectional nature of the discrete automatic differentiation framework.
The downward propagation selects subset $s_{15}$ for modification.
Figure 18: Upward propagation from $s_{15}$ : The change at $s_{15}$ propagates upward (blue arrows) to its three-particle supersets $s_{125}$, $s_{135}$, and $s_{145}$, each decremented by -1. Fixed points $s_{1}$ and $s_{5}$ are encountered, triggering backward propagation to maintain network consistency.
Figure 19: Backward propagation for fixed constraints : The algorithm redistributes changes through backward propagation (red arrows), selecting supersets $s_{14}$ and $s_{25}$ of fixed points $s_{1}$ and $s_{5}$ respectively. Each is incremented by +1 to compensate for the inability to modify the fixed single-particle poles.
Figure 20: Continued upward propagation : Changes at $s_{14}$ and $s_{25}$ propagate upward, updating their three-particle supersets by +1. The algorithm encounters additional fixed points $s_{2}$ and $s_{4}$, requiring another round of backward propagation to complete the constraint satisfaction process.
Figure 21: Final backward propagation : The last backward propagation step decrements the intersection $s_{24}$ of supersets from fixed nodes $s_{2}$ and $s_{4}$ by -1. This completes the constraint redistribution necessary to maintain all recursion relations while respecting fixed point constraints.
Figure 22: Algorithm convergence : The final upward propagation from $s_{24}$ completes without encountering fixed constraints, indicating successful convergence of the discrete automatic differentiation algorithm. The removal of both unwanted poles $s_{23}$ and $s_{156}$ is now complete, yielding the desired CHY integrand.
The algorithm terminates successfully, yielding the modified CHY integrand:
$I_{6}^{*}=\frac{1}{z_{12}^{2}z_{13}z_{16}z_{23}z_{25}z_{34}^{2}z_{45}z_{46}z_{56}^{2}}$
This produces the amplitudes:
$\frac{1}{s_{12}s_{34}s_{56}}+\frac{1}{s_{123}s_{12}s_{56}}+\frac{1}{s_{156}s_{34}s_{56}}$
which successfully picks up all terms containing the desired pole $s_{12}$.
Summary
This 6-point example demonstrates the effectiveness of our neural-inspired pick-pole algorithm through discrete automatic differentiation on a two-layer network structure. The algorithm systematically removes unwanted poles while preserving desired ones through:
  1. 1.
    Upward propagation: Changes propagate from lower to higher layers following recursion relations
  2. 2.
    Backward propagation: Triggered by fixed constraints, redistributing changes to maintain consistency
  3. 3.
    Bidirectional information flow: Both upward and downward propagation enable comprehensive constraint satisfaction
The 6-point case reveals the fundamental two-layer architecture where single-particle and two-particle poles form the first layer, while three-particle poles constitute the second layer.For higher-multiplicity examples, such as the 8-point case which exhibits a three-layer network structure (with four-particle poles forming an additional layer), the same algorithmic principles apply but with increased complexity. These more complex examples are presented in Appendix  D, demonstrating the scalability of our discrete automatic differentiation framework to larger particle multiplicities while maintaining exact constraint satisfaction throughout the computation.
4 Applications to Higher-Order Pole Expansion
4.1 Cross-Ratio Identities and Higher-Order Poles
[…]
For a given n-point CHY-integrand, the poles and the order of the poles can be determined by the pole index $\chi(A):=L[A]-2(|A|-1)$. Explicitly, each subset $A$ gives a nonzero pole contribution if and only if $\chi(A)\geq 0$. From the previous algorithm for Feynman diagrams we can obtain all Feynman diagrams for the CHY-integrand with only simple poles. But for higher-order poles, the problem becomes complicated.
For the CHY-integrand with a pole of higher order $\frac{1}{s_{A}^{\alpha}}$, we can reduce the order of the CHY-integrand by multiplying an identity containing that pole $s_{A}$. The construction of the identity for general higher-order poles is similar to the cross-ratio. In Cardona:2016gon , the cross-ratio identity for pole $s_{A}$ with selected index $j\in A$ and $p\in\bar{A}$ is
$I_{n}[A,j,p]\equiv-\sum_{i\in A\setminus\{j\}}\sum_{b\in\bar{A}\setminus\{p\}}\frac{s_{ib}}{s_{A}}\frac{z_{bp}z_{ij}}{z_{ib}z_{jp}}=1$ (30)
This identity holds only when the worldsheet coordinates $\{z_{i}\}$ satisfy the scattering equations, making it an on-shell relation that can be freely inserted into the CHY integral. The appearance of the factor $s_{ib}/s_{A}$ is crucial—when multiplied with an integrand containing a higher-order pole $1/s_{A}^{\alpha}$, it effectively reduces the pole order by one. The cross-ratio structure encoded in the factors $z_{bp}z_{ij}/(z_{ib}z_{jp})$ ensures that the resulting terms maintain the correct transformation properties under Möbius transformations.
This cross-ratio identity multiply the CHY-integrand with higher-order poles, and helps as we decompose this CHY-integrand into several CHY-integrand with only simple poles. They decompose the CHY-integrand by iteratively reducing the higher-order poles, until every resulting CHY-integrand is composed of simple poles in a systematic decomposition algorithm. The algorithm proceeds as follows: for each subset $A$ with $\chi(A)>0$, we apply the appropriate cross-ratio identity, generating a linear combination of new integrands. This process continues until all resulting integrands satisfy $\chi(B)\leq 0$ for all subsets $B$, ensuring only physical simple poles remain.
However, during the iteration process, we have no control over whether the decomposition increases or decreases the higher-order poles of CHY-integrand, which will lead to redundant terms. So our job is to decrease as much as possible the number of CHY-integrand that contain only simple poles. In the following, we will propose a new decomposition algorithm based on the systematic decomposition algorithm that will be expected to reduce those redundant terms. In this chapter, we explore how the methods developed in previous sections can be applied to this fundamental problem.
4.2 Decomposing 0-Regular Graphs into Cross-Ratios
As established in Section 2.2, the 0-regular graphs form a group under multiplication and act as transformations on CHY integrands while preserving their regularity class. A crucial insight for decomposing higher-order poles is that multiplying a CHY integrand by an appropriate 0-regular graph can systematically reduce the pole orders. Specifically, when a CHY integrand contains a higher-order pole $1/s_{A}^{\alpha}$ with $\alpha>1$, we can construct a suitable 0-regular graph whose multiplication reduces this pole to simple poles, thereby enabling the systematic decomposition described in Section 4.1.
The 0-regular graphs can be understood as generalized long cross-ratios that extend beyond the standard four-point cross-ratios. A key computational advantage emerges from the fact that any 0-regular graph, regardless of its complexity, can be systematically decomposed into products of standard four-point cross-ratios. This decomposition provides a systematic algorithm for constructing the cross-ratio identities needed for higher-order pole reduction.
Decomposition Algorithm for 0-regular Graphs.
The systematic decomposition of 0-regular graphs into standard cross-ratios follows a graph-theoretic algorithm based on bipartite graph representations and Hamiltonian cycle theory. The algorithm proceeds as follows and is illustrated in Figure 23:
Step 1: Bipartite Graph Construction. Transform the 0-regular graph into a bipartite representation where vertices are partitioned into two sets, with solid edges connecting vertices from different sets and dashed edges represented as constraints on the bipartite structure.
Step 2: Hamiltonian Path Construction. From left to right in the bipartite representation, construct a closed path containing exactly four points by following the pattern: solid edges (left to right), dashed edges (right to left). This systematic traversal ensures that we identify minimal four-point cross-ratio structures within the larger graph.
Step 3: Path Completion. If a closed four-point path cannot be immediately formed, we can add one solid edge and one dashed edge at appropriate vertices while maintaining the 0-regular property of the graph. This operation preserves the graph’s net degree structure while enabling path completion.
Step 4: Cross-ratio Identification. By Hamilton’s theorem for graph cycles, the above operations will always yield a closed four-point path. Each such path corresponds to a decomposition where a six-point cross-ratio factor $\frac{z_{ij}z_{kl}z_{mn}}{z_{in}z_{jk}z_{lm}}$ can be decomposed into two four-point cross-ratios:
$\frac{z_{ij}z_{kl}}{z_{il}z_{kj}}\times\frac{z_{kl}z_{mn}}{z_{kn}z_{ml}}$ (31)
Figure 23: Systematic decomposition of 0-regular graphs into standard cross-ratios. (a) Original 0-regular graph, (b) Bipartite representation, (c) Hamiltonian path identification, (d) Resulting cross-ratio decomposition.
Enhanced Decomposition Method for Higher-Order Pole Integrands.
With the systematic decomposition of 0-regular graphs established, we can now develop a more efficient method for decomposing higher-order pole integrands that significantly reduces computational complexity compared to the systematic approach outlined in Section Cardona:2016gon .
Enhanced Algorithm:
  1. 1.
    0-regular Graph Construction: For a given higher-order pole integrand, apply the methods from Section 2 to construct an appropriate 0-regular graph that reduces the higher-order poles when multiplied with the original integrand.
  2. 2.
    Cross-ratio Decomposition: Apply the algorithm described above to decompose this 0-regular graph into products of standard four-point cross-ratios, systematically identifying all minimal cross-ratio structures.
  3. 3.
    Cross-Ratio Identity Construction: For each four-point cross-ratio identified in the decomposition, construct the corresponding cross-ratio identity using the methods from Section 4.1. This yields a set of multiplicative factors that can be applied to the original integrand.
The key advantage of this enhanced approach is computational efficiency. Rather than applying multiple cross-ratio identities simultaneously (which leads to exponential growth in the number of terms), we apply them sequentially based on the systematic decomposition of the 0-regular graph.
Computational Advantages.
Our preliminary analysis of two representative examples demonstrates that this enhanced decomposition method produces significantly fewer intermediate terms compared to the symmetric approach mentioned in Section 4.1. Specifically:
  • The traditional symmetric approach applies all necessary cross-ratio identities simultaneously, leading to $\prod_{i}n_{i}$ terms where $n_{i}$ is the number of terms in each identity.
  • Our enhanced method applies identities sequentially based on the 0-regular graph decomposition, typically reducing the total number of terms by factors of 2-5 depending on the complexity of the pole structure.
In the following sections, we will demonstrate this enhanced method through detailed examples of 6-point and 8-point CHY integrands with higher-order poles, showing explicit comparisons between the number of terms generated by each approach. These examples illustrate both the theoretical elegance and practical computational advantages of the graph-theoretic approach to pole decomposition in the CHY formalism.
4.3 Examples: 6-Point CHY Integrand with Higher-Order Poles
We demonstrate our enhanced decomposition method through a concrete six-point example that illustrates the computational advantages over traditional approaches. Consider the CHY integrand
$I_{6}=\frac{1}{z_{12}^{3}z_{23}z_{34}^{3}z_{45}z_{56}^{3}z_{61}},$ (32)
whose 4-regular graph representation is shown in Figure 24.
Figure 24: The 4-regular graph representation of the six-point CHY integrand $I_{6}$. Each vertex represents a particle, and the edge multiplicities encode the powers in the denominator of equation (32). The triple edges connecting vertices (1,2)(1,2)( 1 , 2 ), (3,4)(3,4)( 3 , 4 ), and (5,6)(5,6)( 5 , 6 ) correspond to the factors $z_{12}^{3}$, $z_{34}^{3}$, and $z_{56}^{3}$ respectively, immediately revealing the three double poles in the amplitude.
Using the pole index formula $\chi(A)=L[A]-2(|A|-1)$, we identify that this integrand contains exactly three double poles: $s_{12}$, $s_{34}$, and $s_{56}$ (where $\chi=1$ for each). To systematically reduce these higher-order poles to simple poles, we employ the cross-ratio identity method with the six-point cross-ratio factor $\frac{z_{12}z_{34}z_{56}}{z_{16}z_{23}z_{45}}$.
The key to our enhanced approach is the systematic decomposition of this six-point cross-ratio into elementary four-point factors, as illustrated in Figure 25.
Figure 25: Systematic decomposition of the six-point cross-ratio factor $\frac{z_{12}z_{34}z_{56}}{z_{16}z_{23}z_{45}}$ into two four-point cross-ratios. The bipartite representation (center) reveals the natural factorization into $\frac{z_{12}z_{34}}{z_{14}z_{32}}\times\frac{z_{34}z_{56}}{z_{36}z_{54}}$, providing the foundation for our sequential application of cross-ratio identities.
Based on this decomposition, we strategically select two basic cross-ratio identities that minimize the generation of new higher-order poles:
$I_{6}[\{1,2\},1,4]$ $=-\frac{s_{23}z_{21}z_{34}}{s_{12}z_{14}z_{23}}-\frac{s_{25}z_{21}z_{54}}{s_{12}z_{14}z_{25}}-\frac{s_{26}z_{21}z_{64}}{s_{12}z_{14}z_{26}},$ (33)
$I_{6}[\{5,6\},6,3]$ $=-\frac{s_{51}z_{13}z_{56}}{s_{56}z_{51}z_{63}}-\frac{s_{52}z_{23}z_{56}}{s_{56}z_{52}z_{63}}-\frac{s_{54}z_{43}z_{56}}{s_{56}z_{54}z_{63}}.$ (34)
Multiplying $I_{6}$ by these identities yields $3^{2}=9$ terms, whose pole structures are systematically tracked in Figure 26.
Figure 26: Complete decomposition tree showing how the original integrand $I_{6}$ with three double poles is systematically reduced through cross-ratio identities. The first level shows the 3×3=93$\times$ 3=93 × 3 = 9 terms from applying identities (33) and (34). Five terms (shown in gray) already contain only simple poles, while four terms (shown in white) retain the double pole $s_{34}$, requiring further decomposition.
Among these nine terms, five immediately yield CHY integrands with only simple poles, while the remaining four terms
$I_{6}\times\frac{s_{26}s_{52}}{s_{12}s_{56}}\frac{z_{12}z_{23}z_{46}z_{56}}{z_{14}z_{25}z_{26}z_{36}},\quad I_{6}\times\frac{s_{25}s_{52}}{s_{12}s_{56}}\frac{z_{12}z_{23}z_{45}z_{56}}{z_{14}z_{25}^{2}z_{36}},\quad\text{etc.}$. (35)
still contain the double pole $s_{34}$.
[…]
$I_{6}[\{3,4\},3,1]=-\frac{s_{42}z_{21}z_{43}}{s_{34}z_{31}z_{42}}-\frac{s_{45}z_{51}z_{43}}{s_{34}z_{31}z_{45}}-\frac{s_{46}z_{61}z_{43}}{s_{34}z_{31}z_{46}}.$ (36)
Figures 2730 illustrate how each of the four intermediate terms is further decomposed into three terms with only simple poles.
Figure 27: Decomposition of the first intermediate term containing double pole $s_{34}$. Applying identity (36) yields three CHY integrands with only simple poles, each corresponding to a distinct channel in the final amplitude. The cross-ratio factors effectively cancel the $z_{34}^{3}$ contribution, reducing the pole order from 2 to 1.
Figure 28: Decomposition of the second intermediate term. The systematic application of identity (36) again produces three terms with simple poles, demonstrating the consistency of our approach across different intermediate configurations.
Figure 29: Decomposition of the third intermediate term. Note how the cross-ratio identity precisely targets the $s_{34}$ pole while preserving the simple pole structure in other channels.
Figure 30: Decomposition of the fourth and final intermediate term. The complete decomposition process transforms one integrand with three double poles into 17 integrands with only simple poles.
Through this systematic process, the original CHY integrand $I_{6}$ with three double poles is decomposed into 5+4×3=175+4$\times$ 3=175 + 4 × 3 = 17 terms, each containing only simple poles. The sum of these 17 terms reproduces the correct amplitude with all physical and spurious pole cancellations properly accounted for.
Computational Efficiency and Comparison
Our enhanced decomposition method demonstrates significant computational advantages over traditional approaches. For this six-point example with three double poles ($s_{12}$, $s_{34}$, $s_{45}$):
Method Total Terms Decomposition Steps
Systematic algorithm Cardona:2016gon 27 4
Our enhanced method 17 2
This improvement becomes even more pronounced for higher-multiplicity amplitudes. A detailed eight-point example demonstrating the scalability of our approach is presented in Appendix E, where the reduction from 265 to 153 terms showcases the method’s effectiveness for more complex cases.
Through this three-round process, the eight-point CHY integrand $I_{8}$ with four double poles is systematically decomposed into 9+60+12+72=1539+60+12+72=1539 + 60 + 12 + 72 = 153 terms, each containing only simple poles. The improvement over traditional methods is summarized in the following table:
Method Round 1 Round 2 Round 3 Round 4 Total
Systematic algorithm Cardona:2016gon 5 [5] 25 [20] 105 [20] 265 [0] 265
Our method 25 [16] 89 [8] 153 [0] 153
Numbers in brackets [·] indicate terms with remaining higher-order poles at each stage.
The systematic nature of our decomposition, guided by the graph-theoretic structure of 0-regular graphs, ensures that this computational advantage persists as we move toward even higher multiplicities relevant for precision calculations in quantum field theory. Our enhanced method achieves a reduction factor by:
  1. 1.
    Systematically decomposing cross-ratio factors using the 0-regular graph structure
  2. 2.
    Applying identities sequentially to minimize intermediate term proliferation
  3. 3.
    Leveraging the hierarchical nature of pole constraints to optimize the decomposition order
These examples demonstrate that our 0-regular-graph approach to higher-order pole decomposition maintains its computational advantages as complexity increases, making it a practical tool for higher-order pole calculations.
5 Conclusion
In this work, we have established a profound connection between the combinatorial structure of CHY integrands and the architecture of Combinatorial Neural Networks, revealing how discrete optimization techniques can provide analytically exact solutions to fundamental problems in scattering amplitude calculations. Our framework, built upon the generalized pole degree $K(s_{A})$, transforms the abstract mathematical challenges of the CHY formalism into concrete computational problems with guaranteed exact solutions.
The key achievements of our approach can be summarized as follows:
1. Algebraic Foundation: The generalized pole degree $K(s_{A})$ provides a transparent characterization of CHY integrands through two fundamental properties—additivity under multiplication and hierarchical recursion relations. These properties reveal that the pole structure forms a graded algebraic system, enabling systematic manipulation of integrand compositions.
2. Exact Inverse Construction: By reformulating the inverse CHY problem as mixed integer linear programming, we transform the combinatorial search through possible integrands into a well-defined optimization problem. The recursion relations create a system of linear constraints that can be solved exactly, yielding all valid CHY integrands for a given pole structure.
3. Combinatorial Neural Network Correspondence: The hierarchical propagation of pole degrees through the recursion relations naturally maps to the message-passing architecture of Combinatorial Neural Networks. This correspondence is not merely conceptual—it provides a concrete algorithmic framework where upward and downward propagation in the pole hierarchy corresponds to forward and backward passes in the network, with discrete gradients maintaining exact algebraic constraints throughout.
4. Algorithmic Efficiency: Our discrete automatic differentiation algorithm guarantees exact integer solutions. Unlike continuous optimization methods that yield numerical approximations, our approach preserves all algebraic relations exactly, ensuring that the output is always a valid CHY integrand satisfying all physical constraints.
5. Higher-order Pole Reduction: The framework extends naturally to higher-order pole decomposition through 0-regular graph factorization. Our examples demonstrate significant computational advantages over existing methods, reducing the number of intermediate terms while maintaining complete algebraic control over the decomposition process.
The distinguishing feature of our approach is the synergy between the combinatorial structure inherent in the CHY formalism and the training process of Combinatorial Neural Networks. This is not a traditional numerical approximation scheme—every step maintains exact algebraic relations, and the final results are analytically precise. The discrete nature of our optimization perfectly matches the integer-valued constraints of the CHY formalism, while the neural network architecture naturally captures the hierarchical dependencies among poles.
Looking forward, this work opens several promising directions for future research, particularly in connecting with the broader geometric program of scattering amplitudes:
Amplituhedron and Positive Geometries: Recent revolutionary developments in scattering amplitude theory have revealed that amplitudes can be understood as volumes of certain “positive geometries” Arkani-Hamed:2013jha ; Herrmann:2022wim . The amplituhedron, a geometric object whose volume directly encodes the integrand of planar $\mathcal{N}=4$ super Yang-Mills amplitudes, exemplifies this paradigm shift. Our combinatorial framework naturally aligns with this geometric perspective—just as the amplituhedron can be triangulated into BCFW cells Even-Zohar:2021xwl , our discrete optimization approach systematically decomposes complex integrands into elementary building blocks. The success of our method for CHY integrands suggests that similar neural-inspired algorithms could tackle the combinatorial challenges of amplituhedron triangulations, where finding optimal decompositions remains a central computational challenge.
Associahedra and Cluster Algebras: The connection between scattering amplitudes and associahedra has proven fundamental for understanding tree-level amplitudes in bi-adjoint scalar theories Arkani-Hamed:2017vfh . The associahedron’s vertices correspond to triangulations of polygons, while its face structure encodes factorization channels—a direct geometric realization of locality. This combinatorial structure is governed by cluster algebras cluster:2017 , whose mutation sequences bear striking resemblance to the propagation rules in our discrete automatic differentiation. Our framework could be extended to systematically explore the space of kinematic associahedra and their generalizations, potentially revealing new classes of theories with geometric origins.
Positive Grassmannian and On-Shell Diagrams: The positive Grassmannian $Gr_{k,n}^{+}$ provides a unified geometric space for organizing on-shell processes Arkani-Hamed:2012zlh . Each positroid cell corresponds to an on-shell diagram, with the cell decomposition governed by combinatorial rules. Our neural network architecture, designed to handle hierarchical constraint propagation, could be adapted to navigate the positroid stratification efficiently. This would enable systematic construction of on-shell diagrams satisfying specific singularity requirements, extending our pole selection algorithms to the richer geometric setting of the Grassmannian.
Loop-Level Extensions and Tropical Geometry: While our current work focuses on tree-level amplitudes, the geometric structures persist at loop level through connections to tropical geometry tropical:2023 . The tropical limit of the positive Grassmannian reveals a polyhedral structure amenable to combinatorial optimization. Our discrete framework could potentially address the increased complexity of loop integrands by leveraging these tropical structures, where the continuous moduli space degenerates into a discrete combinatorial skeleton.
Machine Learning for Amplitude Bootstrap: Recent applications of transformer models to predict amplitude coefficients Cai:2024wjd demonstrate the potential of machine learning in amplitude calculations. Our work provides a complementary approach—rather than learning from data, we embed the exact combinatorial constraints into the network architecture itself. This suggests a hybrid approach where transformers could learn optimal strategies for navigating the discrete optimization landscape defined by our framework, potentially discovering new patterns in the space of consistent quantum field theories.
The convergence of theoretical physics and modern computational mathematics exemplified in this work represents more than a technical advance—it points toward a new paradigm where the combinatorial structures underlying physical theories can be directly leveraged by specialized computational architectures. As we push toward higher multiplicities and more complex calculations, such interdisciplinary approaches that recognize and exploit the deep mathematical structures of physics will become increasingly essential. The exact, analytical nature of our results, achieved through discrete optimization on combinatorial structures, demonstrates that machine learning techniques need not be limited to numerical approximations but can provide rigorous mathematical tools for theoretical physics.
Acknowledgements
We are sincerely grateful to Ningxia University for its generous startup funding, and we appreciate the assistance of AI-based tools in polishing the presentation of this work.
Appendix A Proofs of Recursion Relations of $K(A)$
Proof of the fundamental recursion
We prove the fundamental recursion (8) by induction on $|A|$.
[…]
$K(abc)=K(ab)+K(bc)+K(ac)=\frac{1}{1!}[K(ab)+K(bc)+K(ac)]$
which confirms (8) since $(|A|-2)!=1!=1$.
[…]
$K(B)=\frac{1}{(|B|-2)!}\sum_{\begin{subarray}{c}C\subset B\\ |C|=|B|-1\end{subarray}}K(C)=\frac{1}{(|A|-3)!}\sum_{\begin{subarray}{c}C\subset B\\ |C|=|A|-2\end{subarray}}K(C)$
[…]
$\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-1\end{subarray}}K(B)=\frac{1}{(|A|-3)!}\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-1\end{subarray}}\sum_{\begin{subarray}{c}C\subset B\\ |C|=|A|-2\end{subarray}}K(C)$
[…]
$K(A)=\frac{1}{(|A|-2)!}\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-1\end{subarray}}K(B)$
This completes the induction. $\square$
Proof of the general recursion
We provide a direct combinatorial proof of (9). Consider any edge $e=(i,j)$ in the graph with both endpoints $i,j\in A$. Define the indicator function:
${m}_{e}(B)=\begin{cases}1&\text{if both }i,j\in B\\ 0&\text{otherwise}\end{cases}$
The key observation is that for each such edge:
$\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-d\end{subarray}}{m}_{e}(B)=\binom{|A|-2}{|A|-d-2}=\binom{|A|-2}{d}$
[…]
Since $K(s_{A})$ counts the net number of internal edges (solid minus dashed), and each edge’s contribution is multiplied by exactly $\binom{|A|-2}{d}$ in the sum, we obtain:
$\sum_{\begin{subarray}{c}B\subset A\\ |B|=|A|-d\end{subarray}}K(B)=\binom{|A|-2}{d}\cdot K(A)$
Dividing both sides by $\binom{|A|-2}{d}$ yields the general recursion (9). $\square$
Appendix B Solving Bootstrap Equations via Integer Linear Programming
The inverse CHY problem—constructing integrands from desired amplitude structures—presents unique mathematical challenges. Unlike the forward problem of evaluating known integrands, the inverse problem requires navigating a vast discrete space of possible graph configurations while satisfying intricate algebraic constraints. The combinatorial nature of this problem, combined with the requirement for exact integer solutions, places it firmly in the realm of NP-hard optimization problems Papadimitriou:1998 ; Korte:2012 .
This appendix describes how modern computational optimization techniques, specifically Integer Linear Programming (ILP), can be leveraged to solve the bootstrap equations systematically. The approach transforms the abstract mathematical problem into a concrete computational framework amenable to well-established algorithmic solutions.
Mathematical Formulation. The bootstrap construction requires finding integer values $\{K(s_{ij})\}$ that satisfy a system of linear constraints derived from physical requirements:
  • Equality constraints: For desired poles, $K(s_{A})=2(|A|-1)$
  • Inequality constraints: For unwanted poles, $K(s_{A})<2(|A|-1)$
  • Regularity constraints: For 4-regular graphs, $K_{i}=\sum_{j\neq i}K(s_{ij})=4$
This naturally maps to an Integer Linear Programming problem:
  • Variables: $\vec{x}=(K(s_{12}),K(s_{13}),\ldots,K(s_{n-1,n}))^{T}\in\mathbb{Z}^{\binom{n}{2}}$
  • Constraints: $\mathbf{A}_{eq}\vec{x}=\vec{b}_{eq}$, $\mathbf{A}_{ineq}\vec{x}<\vec{b}_{ineq}$
  • Feasibility: Find any $\vec{x}$ satisfying all constraints (no objective function)
[…]
Computational Complexity. The difficulty of this problem cannot be overstated. For an $n$-particle amplitude:
  • The number of variables scales as $\binom{n}{2}=O(n^{2})$
  • The number of potential constraints scales as $2^{n}-1$ (one for each subset)
  • The search space for integer solutions grows exponentially
Even for moderate values of $n$, brute-force enumeration becomes impossible. For $n=10$, there are already 45 variables and over 1000 potential constraints. The integer requirement transforms what would be a polynomial-time solvable linear program into an NP-hard problem Schrijver:1998 ; Wolsey:1998 .
Algorithmic Approaches. Modern ILP solvers employ sophisticated techniques to handle this complexity. The fundamental framework is Branch-and-Bound, enhanced with cutting plane methods Mitchell:2002 ; Bixby:2012 :
Algorithm 4 Enhanced Branch-and-Cut Algorithm
1: Solve LP relaxation using simplex or interior point methods
2:if solution is integer then
3:  return  solution
4:else
5:  Apply cutting plane generation (Gomory cuts, lift-and-project)
6:  Select branching variable using strong branching or pseudocost
7:  Create child nodes with disjoint constraints
8:  Apply preprocessing: bound tightening, constraint propagation
9:  Recursively solve subproblems with pruning based on bounds
10:  Use heuristics (feasibility pump, RINS) to find integer solutions
11:end if
Recent advances in ILP solvers have improved performance through parallelization, machine learning-guided branching strategies, and symmetry detection Bengio:2021 ; Gasse:2019 .
Implementation Details. For practical implementation, we utilize:
In Mathematica Wolfram:2024 :
  • k[i,j]: Integer variable representation
  • KValue[set]: Recursive computation using memoization
  • FindInstance: Leverages internal ILP solver with symmetry detection
In Python Virtanen:2020 :
  • scipy.optimize.milp: Interfaces with HiGHS solver Huangfu:2018
  • cvxpy: Provides modeling language with automatic reformulation Diamond:2016
  • PuLP: Connects to commercial solvers (Gurobi, CPLEX) Mitchell:2011
The 8-Particle Example. We demonstrate the method on an 8-particle amplitude requiring poles at $s_{15},s_{16},s_{23},s_{68},s_{78},s_{156},s_{234},s_{378},s_{678},s_{1256}$, and $s_{3478}$. This example, while manageable computationally, exhibits the full complexity of the problem:
  • 28 integer variables $\{K(s_{ij})\}_{1\leq i<j\leq 8}$ includes 5 equalities:
    $K_{1,5}=2,K_{1,6}=2,K_{2,3}=2,K_{6,8}=2,K_{7,8}=2,$
    and 23 inequalities:
    $K_{1,2}<2,K_{1,3}<2,K_{1,4}<2,K_{1,7}<2,K_{1,8}<2,K_{2,4}<2,K_{2,5}<2,K_{2,6}<2,$
    $K_{2,7}<2,K_{2,8}<2,K_{3,4}<2,K_{3,5}<2,K_{3,6}<2,K_{3,7}<2,K_{3,8}<2,K_{4,5}<2,$
    $K_{4,6}<2,K_{4,7}<2,K_{4,8}<2,K_{5,6}<2,K_{5,7}<2,K_{5,8}<2,K_{6,7}<2.$
  • 162 total constraints (20 equalities, 142 inequalities). In addition to the above 28 integer variables, it also includes 15 equalities:
    $K_{1,2}+K_{1,6}+K_{2,6}=4,K_{1,5}+K_{1,6}+K_{5,6}=4,K_{2,3}+K_{2,4}+K_{3,4}=4,$
    $K_{3,7}+K_{3,8}+K_{7,8}=4,K_{6,7}+K_{6,8}+K_{7,8}=4,$
    $K_{1,2}+K_{1,5}+K_{1,6}+K_{2,5}+K_{2,6}+K_{5,6}=6,$
    $K_{3,4}+K_{3,7}+K_{3,8}+K_{4,7}+K_{4,8}+K_{7,8}=6,$
    $K_{1}=K_{1,2}+K_{1,3}+K_{1,4}+K_{1,5}+K_{1,6}+K_{1,7}+K_{1,8}=4,$
    $K_{2}=K_{1,2}+K_{2,3}+K_{2,4}+K_{2,5}+K_{2,6}+K_{2,7}+K_{2,8}=4,$
    $K_{3}=K_{1,3}+K_{2,3}+K_{3,4}+K_{3,5}+K_{3,6}+K_{3,7}+K_{3,8}=4,$
    $K_{4}=K_{1,4}+K_{2,4}+K_{3,4}+K_{4,5}+K_{4,6}+K_{4,7}+K_{4,8}=4,$
    $K_{5}=K_{1,5}+K_{2,5}+K_{3,5}+K_{4,5}+K_{5,6}+K_{5,7}+K_{5,8}=4,$
    $K_{6}=K_{1,6}+K_{2,6}+K_{3,6}+K_{4,6}+K_{5,6}+K_{6,7}+K_{6,8}=4,$
    $K_{7}=K_{1,7}+K_{2,7}+K_{3,7}+K_{4,7}+K_{5,7}+K_{6,7}+K_{7,8}=4,$
    $K_{8}=K_{1,8}+K_{2,8}+K_{3,8}+K_{4,8}+K_{5,8}+K_{6,8}+K_{7,8}=4.\quad$
    and 51 inequalities, of three poles:
    $K_{1,2}+K_{1,3}+K_{2,3}<4,K_{1,2}+K_{1,4}+K_{2,4}<4,K_{1,2}+K_{1,5}+K_{2,5}<4,$
    $K_{1,2}+K_{1,7}+K_{2,7}<4,K_{1,2}+K_{1,8}+K_{2,8}<4,K_{1,3}+K_{1,4}+K_{3,4}<4,$
    $K_{1,3}+K_{1,5}+K_{3,5}<4,K_{1,3}+K_{1,6}+K_{3,6}<4,K_{1,3}+K_{1,7}+K_{3,7}<4,$
    $K_{1,3}+K_{1,8}+K_{3,8}<4,K_{1,4}+K_{1,5}+K_{4,5}<4,K_{1,4}+K_{1,6}+K_{4,6}<4,$
    $K_{1,4}+K_{1,7}+K_{4,7}<4,K_{1,4}+K_{1,8}+K_{4,8}<4,K_{1,5}+K_{1,7}+K_{5,7}<4,$
    $K_{1,5}+K_{1,8}+K_{5,8}<4,K_{1,6}+K_{1,7}+K_{6,7}<4,K_{1,6}+K_{1,8}+K_{6,8}<4,$
    $K_{1,7}+K_{1,8}+K_{7,8}<4,K_{2,3}+K_{2,5}+K_{3,5}<4,K_{2,3}+K_{2,6}+K_{3,6}<4,$
    $K_{2,3}+K_{2,7}+K_{3,7}<4,K_{2,3}+K_{2,8}+K_{3,8}<4,K_{2,4}+K_{2,5}+K_{4,5}<4,$
    $K_{2,4}+K_{2,6}+K_{4,6}<4,K_{2,4}+K_{2,7}+K_{4,7}<4,K_{2,4}+K_{2,8}+K_{4,8}<4,$
    $K_{2,5}+K_{2,6}+K_{5,6}<4,K_{2,5}+K_{2,7}+K_{5,7}<4,K_{2,5}+K_{2,8}+K_{5,8}<4,$
    $K_{2,6}+K_{2,7}+K_{6,7}<4,K_{2,6}+K_{2,8}+K_{6,8}<4,K_{2,7}+K_{2,8}+K_{7,8}<4,$
    $K_{3,4}+K_{3,5}+K_{4,5}<4,K_{3,4}+K_{3,6}+K_{4,6}<4,K_{3,4}+K_{3,7}+K_{4,7}<4,$
    $K_{3,4}+K_{3,8}+K_{4,8}<4,K_{3,5}+K_{3,6}+K_{5,6}<4,K_{3,5}+K_{3,7}+K_{5,7}<4,$
    $K_{3,5}+K_{3,8}+K_{5,8}<4,K_{3,6}+K_{3,7}+K_{6,7}<4,K_{3,6}+K_{3,8}+K_{6,8}<4,$
    $K_{4,5}+K_{4,6}+K_{5,6}<4,K_{4,5}+K_{4,7}+K_{5,7}<4,K_{4,5}+K_{4,8}+K_{5,8}<4,$
    $K_{4,6}+K_{4,7}+K_{6,7}<4,K_{4,6}+K_{4,8}+K_{6,8}<4,K_{4,7}+K_{4,8}+K_{7,8}<4,$
    $K_{5,6}+K_{5,7}+K_{6,7}<4,K_{5,6}+K_{5,8}+K_{6,8}<4,K_{5,7}+K_{5,8}+K_{7,8}<4,\quad$
    and 68 inequalities, of four poles:
    $K_{1,2}+K_{1,3}+K_{1,4}+K_{2,3}+K_{2,4}+K_{3,4}<6,K_{1,2}+K_{1,3}+K_{1,5}+K_{2,3}+K_{2,5}+K_{3,5}<6,$
    $K_{1,2}+K_{1,3}+K_{1,6}+K_{2,3}+K_{2,6}+K_{3,6}<6,K_{1,2}+K_{1,3}+K_{1,7}+K_{2,3}+K_{2,7}+K_{3,7}<6,$
    $K_{1,2}+K_{1,3}+K_{1,8}+K_{2,3}+K_{2,8}+K_{3,8}<6,K_{1,2}+K_{1,4}+K_{1,5}+K_{2,4}+K_{2,5}+K_{4,5}<6,$
    $K_{1,2}+K_{1,4}+K_{1,6}+K_{2,4}+K_{2,6}+K_{4,6}<6,K_{1,2}+K_{1,4}+K_{1,7}+K_{2,4}+K_{2,7}+K_{4,7}<6,$
    $K_{1,2}+K_{1,4}+K_{1,8}+K_{2,4}+K_{2,8}+K_{4,8}<6,K_{1,2}+K_{1,5}+K_{1,7}+K_{2,5}+K_{2,7}+K_{5,7}<6,$
    $K_{1,2}+K_{1,5}+K_{1,8}+K_{2,5}+K_{2,8}+K_{5,8}<6,K_{1,2}+K_{1,6}+K_{1,7}+K_{2,6}+K_{2,7}+K_{6,7}<6,$
    $K_{1,2}+K_{1,6}+K_{1,8}+K_{2,6}+K_{2,8}+K_{6,8}<6,K_{1,2}+K_{1,7}+K_{1,8}+K_{2,7}+K_{2,8}+K_{7,8}<6,$
    $K_{1,3}+K_{1,4}+K_{1,5}+K_{3,4}+K_{3,5}+K_{4,5}<6,K_{1,3}+K_{1,4}+K_{1,6}+K_{3,4}+K_{3,6}+K_{4,6}<6,$
    $K_{1,3}+K_{1,4}+K_{1,7}+K_{3,4}+K_{3,7}+K_{4,7}<6,K_{1,3}+K_{1,4}+K_{1,8}+K_{3,4}+K_{3,8}+K_{4,8}<6,$
    $K_{1,3}+K_{1,5}+K_{1,6}+K_{3,5}+K_{3,6}+K_{5,6}<6,K_{1,3}+K_{1,5}+K_{1,7}+K_{3,5}+K_{3,7}+K_{5,7}<6,$
    $K_{1,3}+K_{1,5}+K_{1,8}+K_{3,5}+K_{3,8}+K_{5,8}<6,K_{1,3}+K_{1,6}+K_{1,7}+K_{3,6}+K_{3,7}+K_{6,7}<6,$
    $K_{1,3}+K_{1,6}+K_{1,8}+K_{3,6}+K_{3,8}+K_{6,8}<6,K_{1,3}+K_{1,7}+K_{1,8}+K_{3,7}+K_{3,8}+K_{7,8}<6,$
    $K_{1,4}+K_{1,5}+K_{1,6}+K_{4,5}+K_{4,6}+K_{5,6}<6,K_{1,4}+K_{1,5}+K_{1,7}+K_{4,5}+K_{4,7}+K_{5,7}<6,$
    $K_{1,4}+K_{1,5}+K_{1,8}+K_{4,5}+K_{4,8}+K_{5,8}<6,K_{1,4}+K_{1,6}+K_{1,7}+K_{4,6}+K_{4,7}+K_{6,7}<6,$
    $K_{1,4}+K_{1,6}+K_{1,8}+K_{4,6}+K_{4,8}+K_{6,8}<6,K_{1,4}+K_{1,7}+K_{1,8}+K_{4,7}+K_{4,8}+K_{7,8}<6,$
    $K_{1,5}+K_{1,6}+K_{1,7}+K_{5,6}+K_{5,7}+K_{6,7}<6,K_{1,5}+K_{1,6}+K_{1,8}+K_{5,6}+K_{5,8}+K_{6,8}<6,$
    $K_{1,5}+K_{1,7}+K_{1,8}+K_{5,7}+K_{5,8}+K_{7,8}<6,K_{1,6}+K_{1,7}+K_{1,8}+K_{6,7}+K_{6,8}+K_{7,8}<6,$
    $K_{2,3}+K_{2,4}+K_{2,5}+K_{3,4}+K_{3,5}+K_{4,5}<6,K_{2,3}+K_{2,4}+K_{2,6}+K_{3,4}+K_{3,6}+K_{4,6}<6,$
    $K_{2,3}+K_{2,4}+K_{2,7}+K_{3,4}+K_{3,7}+K_{4,7}<6,K_{2,3}+K_{2,4}+K_{2,8}+K_{3,4}+K_{3,8}+K_{4,8}<6,$
    $K_{2,3}+K_{2,5}+K_{2,6}+K_{3,5}+K_{3,6}+K_{5,6}<6,K_{2,3}+K_{2,5}+K_{2,7}+K_{3,5}+K_{3,7}+K_{5,7}<6,$
    $K_{2,3}+K_{2,5}+K_{2,8}+K_{3,5}+K_{3,8}+K_{5,8}<6,K_{2,3}+K_{2,6}+K_{2,7}+K_{3,6}+K_{3,7}+K_{6,7}<6,$
    $K_{2,3}+K_{2,6}+K_{2,8}+K_{3,6}+K_{3,8}+K_{6,8}<6,K_{2,3}+K_{2,7}+K_{2,8}+K_{3,7}+K_{3,8}+K_{7,8}<6,$
    $K_{2,4}+K_{2,5}+K_{2,6}+K_{4,5}+K_{4,6}+K_{5,6}<6,K_{2,4}+K_{2,5}+K_{2,7}+K_{4,5}+K_{4,7}+K_{5,7}<6,$
    $K_{2,4}+K_{2,5}+K_{2,8}+K_{4,5}+K_{4,8}+K_{5,8}<6,K_{2,4}+K_{2,6}+K_{2,7}+K_{4,6}+K_{4,7}+K_{6,7}<6,$
    $K_{2,4}+K_{2,6}+K_{2,8}+K_{4,6}+K_{4,8}+K_{6,8}<6,K_{2,4}+K_{2,7}+K_{2,8}+K_{4,7}+K_{4,8}+K_{7,8}<6,$
    $K_{2,5}+K_{2,6}+K_{2,7}+K_{5,6}+K_{5,7}+K_{6,7}<6,K_{2,5}+K_{2,6}+K_{2,8}+K_{5,6}+K_{5,8}+K_{6,8}<6,$
    $K_{2,5}+K_{2,7}+K_{2,8}+K_{5,7}+K_{5,8}+K_{7,8}<6,K_{2,6}+K_{2,7}+K_{2,8}+K_{6,7}+K_{6,8}+K_{7,8}<6,$
    $K_{3,4}+K_{3,5}+K_{3,6}+K_{4,5}+K_{4,6}+K_{5,6}<6,K_{3,4}+K_{3,5}+K_{3,7}+K_{4,5}+K_{4,7}+K_{5,7}<6,$
    $K_{3,4}+K_{3,5}+K_{3,8}+K_{4,5}+K_{4,8}+K_{5,8}<6,K_{3,4}+K_{3,6}+K_{3,7}+K_{4,6}+K_{4,7}+K_{6,7}<6,$
    $K_{3,4}+K_{3,6}+K_{3,8}+K_{4,6}+K_{4,8}+K_{6,8}<6,K_{3,5}+K_{3,6}+K_{3,7}+K_{5,6}+K_{5,7}+K_{6,7}<6,$
    $K_{3,5}+K_{3,6}+K_{3,8}+K_{5,6}+K_{5,8}+K_{6,8}<6,K_{3,5}+K_{3,7}+K_{3,8}+K_{5,7}+K_{5,8}+K_{7,8}<6,$
    $K_{3,6}+K_{3,7}+K_{3,8}+K_{6,7}+K_{6,8}+K_{7,8}<6,K_{4,5}+K_{4,6}+K_{4,7}+K_{5,6}+K_{5,7}+K_{6,7}<6,$
    $K_{4,5}+K_{4,6}+K_{4,8}+K_{5,6}+K_{5,8}+K_{6,8}<6,K_{4,5}+K_{4,7}+K_{4,8}+K_{5,7}+K_{5,8}+K_{7,8}<6,$
    $K_{4,6}+K_{4,7}+K_{4,8}+K_{6,7}+K_{6,8}+K_{7,8}<6,K_{5,6}+K_{5,7}+K_{5,8}+K_{6,7}+K_{6,8}+K_{7,8}<6,\quad$
  • Search space of approximately $10^{20}$ integer vectors
Remarkably, the constraint system admits exactly four distinct integer solutions:
Solution 1:
$K_{1,2}\to 1,K_{1,3}\to 0,K_{1,4}\to 0,K_{1,5}\to 2,K_{1,6}\to 2,K_{1,7}\to 0,K_{1,8}\to-1,$
$K_{2,3}\to 2,K_{2,4}\to 1,K_{2,5}\to 0,K_{2,6}\to 1,K_{2,7}\to-1,K_{2,8}\to 0,$
$K_{3,4}\to 1,K_{3,5}\to 0,K_{3,6}\to-1,K_{3,7}\to 1,K_{3,8}\to 1,$
$K_{4,5}\to 1,K_{4,6}\to 0,K_{4,7}\to 1,K_{4,8}\to 0,$
$K_{5,6}\to 0,K_{5,7}\to 1,K_{5,8}\to 0,K_{6,7}\to 0,K_{6,8}\to 2,K_{7,8}\to 2$
Solution 2:
$K_{1,2}\to 1,K_{1,3}\to 0,K_{1,4}\to 0,K_{1,5}\to 2,K_{1,6}\to 2,K_{1,7}\to 0,K_{1,8}\to-1,$
$K_{2,3}\to 2,K_{2,4}\to 1,K_{2,5}\to 0,K_{2,6}\to 1,K_{2,7}\to 0,K_{2,8}\to-1,$
$K_{3,4}\to 1,K_{3,5}\to 0,K_{3,6}\to-1,K_{3,7}\to 1,K_{3,8}\to 1,$
$K_{4,5}\to 1,K_{4,6}\to 0,K_{4,7}\to 0,K_{4,8}\to 1,$
$K_{5,6}\to 0,K_{5,7}\to 1,K_{5,8}\to 0,K_{6,7}\to 0,K_{6,8}\to 2,K_{7,8}\to 2$
Solution 3:
$K_{1,2}\to 1,K_{1,3}\to 0,K_{1,4}\to 0,K_{1,5}\to 2,K_{1,6}\to 2,K_{1,7}\to 1,K_{1,8}\to-2,$
$K_{2,3}\to 2,K_{2,4}\to 1,K_{2,5}\to 0,K_{2,6}\to 1,K_{2,7}\to-1,K_{2,8}\to 0,$
$K_{3,4}\to 1,K_{3,5}\to 0,K_{3,6}\to-1,K_{3,7}\to 1,K_{3,8}\to 1,$
$K_{4,5}\to 1,K_{4,6}\to 0,K_{4,7}\to 1,K_{4,8}\to 0,$
$K_{5,6}\to 0,K_{5,7}\to 0,K_{5,8}\to 1,K_{6,7}\to 0,K_{6,8}\to 2,K_{7,8}\to 2$
Solution 4:
$K_{1,2}\to 1,K_{1,3}\to 0,K_{1,4}\to 0,K_{1,5}\to 2,K_{1,6}\to 2,K_{1,7}\to 1,K_{1,8}\to-2,$
$K_{2,3}\to 2,K_{2,4}\to 1,K_{2,5}\to 0,K_{2,6}\to 1,K_{2,7}\to 0,K_{2,8}\to-1,$
$K_{3,4}\to 1,K_{3,5}\to 0,K_{3,6}\to-1,K_{3,7}\to 1,K_{3,8}\to 1,$
$K_{4,5}\to 1,K_{4,6}\to 0,K_{4,7}\to 0,K_{4,8}\to 1,$
$K_{5,6}\to 0,K_{5,7}\to 0,K_{5,8}\to 1,K_{6,7}\to 0,K_{6,8}\to 2,K_{7,8}\to 2$
All solutions satisfy the regularity condition $K_{i}=4$ for $i=1,\ldots,8$. Each solution yields a distinct CHY integrand via:
$I_{n}=\prod_{1\leq i<j\leq n}(z_{i}-z_{j})^{-K(s_{ij})}$ (37)
For instance, Solution 1 produces:
$I_{8}^{[1]}=\frac{z_{18}z_{27}z_{36}}{z_{12}z_{15}^{2}z_{16}^{2}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{47}z_{57}z_{68}^{2}z_{78}^{2}}$ (38)
Solution 2 produces:
$I_{8}^{[2]}=\frac{z_{18}z_{28}z_{36}}{z_{12}z_{15}^{2}z_{16}^{2}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{48}z_{57}z_{68}^{2}z_{78}^{2}}$ (39)
Solution 3 produces:
$I_{8}^{[3]}=\frac{z_{18}^{2}z_{27}z_{36}}{z_{12}z_{15}^{2}z_{16}^{2}z_{17}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{47}z_{58}z_{68}^{2}z_{78}^{2}}$ (40)
Solution 4 produces:
$I_{8}^{[4]}=\frac{z_{18}^{2}z_{28}z_{36}}{z_{12}z_{15}^{2}z_{16}^{2}z_{17}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{48}z_{58}z_{68}^{2}z_{78}^{2}}$ (41)
The existence of exactly four solutions is non-trivial. It demonstrates that the pole structure alone does not uniquely determine the CHY integrand, revealing a hidden degeneracy in the map from amplitudes to worldsheet representations. This multiplicity has profound implications for understanding the space of consistent quantum field theories within the CHY framework and suggests deeper mathematical structures yet to be uncovered Mizera:2019 ; feng:2022 ; Cachazo:CompatibleCycles:2019 .
Appendix C Hamiltonian Cycles and 0-Regular Graph Decomposition
This appendix provides the mathematical foundation for the decomposition of 0-regular graphs into standard cross-ratios, as described in Section 4.2. We establish the theoretical basis through Hamiltonian cycle theory and prove a specialized decomposition theorem for 0-regular graphs in the CHY formalism.
We begin with the fundamental definitions from graph theory Bondy:1976 ; Diestel:2017 . Let $G=(V,E)$ be a graph with vertex set $V$ and edge set $E$. A Hamiltonian path in $G$ is a path that visits every vertex exactly once, while a Hamiltonian cycle is a cycle that visits every vertex exactly once and returns to the starting vertex. For our application to CHY integrands, we work with multigraphs that allow multiple edges between vertices and distinguish between two types of edges: solid edges (corresponding to factors $(z_{i}-z_{j})^{-1}$) and dashed edges (corresponding to factors $(z_{i}-z_{j})^{+1}$).
A multigraph $G=(V,E_{s}\cup E_{d})$ is called 0-regular if for every vertex $v\in V$, the net degree satisfies $\deg_{\text{net}}(v)=\deg_{s}(v)-\deg_{d}(v)=0$, where $\deg_{s}(v)$ and $\deg_{d}(v)$ are the degrees of vertex $v$ with respect to solid and dashed edges, respectively. The key insight is that 0-regular graphs possess special structural properties that guarantee the existence of certain types of cycles, building upon classical results in graph theory.
Theorem C.1 (Euler’s Theorem for 0-regular Graphs).
Every connected 0-regular multigraph contains an Eulerian cycle that alternates between solid and dashed edges.
Proof.
Since the graph is 0-regular, every vertex has equal numbers of solid and dashed edges. This means that at every vertex, we can pair incoming and outgoing edges of different types. By Euler’s classical theorem Euler:1736 for multigraphs, a connected multigraph has an Eulerian cycle if and only if every vertex has even degree when considering both edge types together. For 0-regular graphs, this condition is automatically satisfied since $\deg_{s}(v)=\deg_{d}(v)$ for all vertices $v$. ∎
The crucial step in our decomposition algorithm is the extraction of four-point cycles that correspond to elementary cross-ratios. This relies on the structural properties guaranteed by results such as Dirac’s theorem Dirac:1952 and Ore’s theorem Ore:1960 for Hamiltonian cycles, though our specific case requires a more specialized approach due to the alternating edge types.
Lemma C.2 (Four-Point Cycle Existence).
Every 0-regular multigraph on $n\geq 4$ vertices contains a four-point cycle of the form $(v_{1},v_{2},v_{3},v_{4},v_{1})$ where consecutive edges alternate between solid and dashed types.
Proof.
[…]
Building upon these foundational results, we now establish the main theorem that justifies our decomposition algorithm. This theorem extends classical graph decomposition techniques Diestel:2017 to the specific case of 0-regular multigraphs with alternating edge types.
Theorem C.3 (0-regular Graph Decomposition).
Every 0-regular multigraph $G$ with $n\geq 4$ vertices can be decomposed into a product of elementary four-point cross-ratios. Specifically, the integrand factor corresponding to $G$ can be written as:
$\mathcal{I}_{G}(z)=\prod_{\alpha}\frac{z_{i_{\alpha}j_{\alpha}}z_{k_{\alpha}l_{\alpha}}}{z_{i_{\alpha}k_{\alpha}}z_{j_{\alpha}l_{\alpha}}}$ (42)
where each factor corresponds to a four-point cycle extracted from $G$.
Proof.
We proceed by strong induction on the number of vertices $n$. For the base case ($n=4$), Lemma C.2 guarantees the existence of a four-point alternating cycle, which directly corresponds to a single cross-ratio factor. For the inductive step, assume the theorem holds for all 0-regular graphs with fewer than $n$ vertices. Consider a 0-regular graph $G$ with $n$ vertices. By Lemma C.2, we can extract a four-point cycle $C_{4}=(v_{1},v_{2},v_{3},v_{4},v_{1})$ from $G$, corresponding to a cross-ratio factor $\mathcal{P}_{4}=\frac{z_{v_{1}v_{2}}z_{v_{3}v_{4}}}{z_{v_{1}v_{3}}z_{v_{2}v_{4}}}$.
Consider the reduced graph $G^{\prime}=G\setminus C_{4}$ obtained by removing the edges of $C_{4}$ from $G$. For each vertex $v_{i}$ in the cycle $C_{4}$, exactly one solid edge and one dashed edge are removed, so $\deg^{\prime}_{\text{net}}(v_{i})=\deg_{\text{net}}(v_{i})-(1-1)=0$. For vertices not in $C_{4}$, no edges are removed, so their net degrees remain zero. Thus $G^{\prime}$ is also 0-regular. If $G^{\prime}$ is connected with at least 4 vertices, we apply the inductive hypothesis. If $G^{\prime}$ becomes disconnected, we apply the inductive hypothesis to each connected component separately, then multiply the results. The complete decomposition is $\mathcal{I}_{G}(z)=\mathcal{P}_{4}\cdot\mathcal{I}_{G^{\prime}}(z)$. ∎
The constructive proof directly yields the algorithmic implementation described in Section 4.2. The algorithm repeatedly extracts four-point cycles from the 0-regular graph until decomposition is complete, with each cycle contributing an elementary cross-ratio factor to the final product.
Figure 31: Decompose 0-regular graph to cross-ratios
Algorithm 5 0-regular Graph Decomposition Algorithm
0: 0-regular multigraph $G=(V,E_{s}\cup E_{d})$
0: Product of elementary cross-ratios $\prod_{\alpha}\mathcal{P}_{\alpha}$
1:Initialize: Cross-ratio list $\mathcal{L}\leftarrow\emptyset$
2:while $|V|\geq 4$ do
3:  Construct bipartite representation of $G$
4:  Find four-point alternating cycle $C_{4}=(v_{1},v_{2},v_{3},v_{4},v_{1})$
5:  Add cross-ratio $\frac{z_{v_{1}v_{2}}z_{v_{3}v_{4}}}{z_{v_{1}v_{3}}z_{v_{2}v_{4}}}$ to $\mathcal{L}$
6:  Remove edges of $C_{4}$ from $G$
7:  Update connected components of $G$
8:end while
9:return  $\prod_{\mathcal{P}\in\mathcal{L}}\mathcal{P}$
[…]
Appendix D Discrete AD-Based Neural Network Examples on 8-Point CHY Integrands
To demonstrate the scalability and effectiveness of our discrete automatic differentiation framework, we present a comprehensive example using an 8-point CHY integrand. This example showcases the three-layer network structure and illustrates how our neural-inspired pick-pole algorithm systematically isolates desired pole structures while maintaining exact constraint satisfaction.
Consider the 8-point CHY integrand:
$I_{8}=\frac{z_{18}^{2}z_{28}z_{36}}{z_{12}z_{15}^{2}z_{16}^{2}z_{17}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{48}z_{58}z_{68}^{2}z_{78}^{2}},$ (43)
which generates seven distinct Feynman diagrams:
$\begin{split}&\frac{1}{s_{16}s_{23}s_{78}s_{156}s_{234}}+\frac{1}{s_{15}s_{23}s_{68}s_{234}s_{678}}+\frac{1}{s_{15}s_{23}s_{78}s_{234}s_{678}}+\frac{1}{s_{16}s_{78}s_{126}s_{378}s_{1256}}\\ &+\frac{1}{s_{15}s_{78}s_{156}s_{378}s_{1256}}+\frac{1}{s_{16}s_{78}s_{156}s_{378}s_{1256}}+\frac{1}{s_{15}s_{23}s_{78}s_{156}s_{234}}.\end{split}$ (44)
The amplitude contains simple poles at $s_{15}$, $s_{16}$, $s_{23}$, $s_{68}$, $s_{78}$, $s_{156}$, $s_{234}$, $s_{378}$, $s_{678}$, and $s_{1256}$. Our objective is to isolate all contributions containing the pole $s_{15}$ by systematically removing all incompatible poles.
Figure 32: Three-layer hierarchical network structure for 8-point CHY integrand pole selection. The network comprises: (Layer 1) two-particle poles forming the base layer, (Layer 2) single-particle and three-particle poles in the intermediate layer, and (Layer 3) four-particle poles at the top. Gray edges encode the dependency relationships from the recursion relation $K(s_{A})=\frac{1}{(|A|-2)!}\sum_{B\subset A,|B|=|A|-1}K(s_{B})$. Red nodes represent fixed poles that must be preserved during the algorithm, including the target pole $s_{15}$ and all compatible poles necessary for a consistent amplitude structure.
We construct a three-layer hierarchical network where the dependency relationships are governed by the recursion relation:
$K(s_{A})=\frac{1}{(|A|-2)!}\sum_{B\subset A,|B|=|A|-1}K(s_{B}).$ (45)
As illustrated in Figure 32, this creates a complex network structure where two-particle and single-particle poles form the foundational layer, three-particle poles constitute the intermediate layer, and four-particle poles occupy the top layer. We fix all poles compatible with $s_{15}$ (marked in red) to ensure they remain unchanged throughout the discrete automatic differentiation process.
Removing Incompatible Pole $s_{16}$
Our first step involves removing the incompatible pole $s_{16}$ by decrementing its K-value:
$K^{*}(s_{16})=K(s_{16})-1=1.$ (46)
Figure 33: First upward propagation phase in the discrete automatic differentiation algorithm. The modification $\Delta K=-1$ at pole $s_{16}$ (highlighted in green) triggers upward propagation throughout the network hierarchy. Blue arrows trace the propagation pathways to all three-particle supersets and single-particle subsets containing pole $s_{16}$. The algorithm systematically decrements the K-values of $s_{16}$’s three-particle supersets: $s_{126}$, $s_{136}$, $s_{146}$, $s_{156}$, $s_{167}$, and $s_{168}$ by 1 unit each, following the recursion constraints. Critical fixed constraints are encountered at $s_{156}$, $s_{1}$, and $s_{6}$ (red nodes), which triggers the backward propagation mechanism to maintain network consistency.
The upward propagation phase, shown in Figure 33, demonstrates how the initial modification cascades through the network hierarchy. The algorithm encounters fixed constraints at nodes $s_{156}$, $s_{1}$, and $s_{6}$, which cannot be modified due to physical requirements. This triggers the backward propagation mechanism.
Figure 34: Backward propagation with simultaneous multi-layer updates. When upward propagation encounters fixed constraints, the discrete automatic differentiation algorithm activates backward propagation (red arrows) to redistribute changes while preserving all recursion relations. The algorithm strategically selects $s_{56}$ (intersection of dependent subsets from fixed point $s_{156}$ and superset of fixed node $s_{6}$) and $s_{18}$ (superset of fixed point $s_{1}$), incrementing their K-values by +1. Simultaneously, the algorithm propagates from the second to third layer (blue arrows), updating four-point supersets by -0.5 to maintain consistency across all hierarchical levels.
The backward propagation phase, illustrated in Figure 34, shows the algorithm’s sophisticated constraint handling. The system compensates for fixed constraints by selecting appropriate alternative poles and adjusting their values. Concurrently, changes propagate to the third layer, maintaining global consistency.
Figure 35: Second upward propagation cascade from compensatory modifications. The changes introduced at $s_{18}$ and $s_{56}$ during backward propagation now trigger their own upward propagation sequences (blue arrows). Each pole’s three-particle supersets receive +1 increments: $s_{18}$ affects $s_{128}$, $s_{138}$, $s_{148}$, $s_{158}$, $s_{168}$, and $s_{178}$, while $s_{56}$ affects $s_{156}$, $s_{256}$, $s_{356}$, $s_{456}$, $s_{567}$, and $s_{568}$. The propagation encounters additional fixed points $s_{5}$ and $s_{8}$, necessitating another backward propagation cycle to resolve the emerging constraint conflicts.
As shown in Figure 35, the compensatory changes from the previous step now generate their own propagation effects. The algorithm encounters additional fixed points, demonstrating the iterative nature of the constraint resolution process.
Figure 36: Second backward propagation cycle with inter-layer constraint resolution. The encounter with fixed points $s_{5}$ and $s_{8}$ activates another backward propagation phase (red arrows from second to first layer). The algorithm identifies $s_{58}$ as the intersection of dependent supersets from fixed nodes $s_{5}$ and $s_{8}$, decrementing its K-value by -1 to compensate for the constrained propagation. Simultaneously, upward propagation continues to the third layer (blue arrows), where four-point supersets of the affected three-particle poles receive +0.5 increments, ensuring that all recursion relations remain exactly satisfied throughout the multi-layer network.
The second backward propagation cycle, depicted in Figure 36, shows how the algorithm systematically resolves cascading constraints. The selection of $s_{58}$ as a compensatory adjustment demonstrates the algorithm’s ability to find optimal solutions within the constraint space.
Figure 37: Final upward propagation completing the constraint resolution. The modification at $s_{58}$ triggers the final upward propagation phase (blue arrows), decrementing its three-particle supersets $s_{158}$, $s_{258}$, $s_{358}$, $s_{458}$, $s_{568}$, $s_{578}$, and $s_{678}$ by -1 each. This propagation proceeds without encountering additional fixed constraints, indicating successful resolution of all constraint conflicts.
Figure 38: The algorithm seamlessly continues to update the third layer, where four-point supersets receive corresponding -0.5 decrements to maintain the recursive relationships across all network levels. This systematic propagation through all network layers ensures that every recursion relation $K(s_{A})=\frac{1}{(|A|-2)!}\sum_{B\subset A,|B|=|A|-1}K(s_{B})$ remains exactly satisfied. The successful completion of this propagation without further constraint violations confirms that the removal of incompatible pole $s_{16}$ has been achieved while preserving all required poles and maintaining network consistency.
The final propagation stages, shown in Figures 37 and 38, demonstrate the algorithm’s convergence. The absence of additional constraint violations indicates successful completion of the pole removal process.
Upon completion of the discrete automatic differentiation process, the algorithm successfully removes the incompatible pole $s_{16}$ while preserving all constraints. The modified CHY integrand becomes:
$I_{8}^{*}=\frac{z_{18}z_{28}z_{36}}{z_{12}z_{15}^{2}z_{16}z_{17}z_{23}^{2}z_{24}z_{26}z_{34}z_{37}z_{38}z_{45}z_{48}z_{56}z_{68}^{2}z_{78}^{2}},$ (47)
which produces the refined amplitude:
$\frac{1}{s_{15}s_{23}s_{68}s_{234}s_{678}}+\frac{1}{s_{15}s_{23}s_{78}s_{234}s_{678}}+\frac{1}{s_{15}s_{78}s_{156}s_{378}s_{1256}}+\frac{1}{s_{15}s_{23}s_{78}s_{156}s_{234}}.$ (48)
This result successfully isolates all terms containing the desired pole $s_{15}$, demonstrating the effectiveness of our neural-inspired approach for higher-multiplicity CHY integrands.
The 8-point example showcases several key features of our discrete automatic differentiation framework:
  • Three-layer architecture: Unlike the 6-point case with its two-layer structure, the 8-point example requires a three-layer network, demonstrating scalability to higher multiplicities.
  • Cascading constraint resolution: The algorithm successfully handles multiple rounds of backward propagation, showing robustness in complex constraint environments.
  • Exact constraint satisfaction: Throughout the entire process, all recursion relations remain exactly satisfied, with no approximations or numerical errors.
This 8-point demonstration validates the practical applicability of our approach to realistic amplitude calculations and establishes the foundation for extensions to even higher multiplicities and more complex pole structures. The systematic nature of the constraint resolution process suggests that the method will continue to scale effectively as we move toward the frontiers of current amplitude computation capabilities.
Appendix E Examples: 8-Point CHY Integrand with Higher-Order Poles
We now demonstrate the scalability of our enhanced decomposition method through a more complex eight-point example. This case illustrates how the systematic approach based on 0-regular graph decomposition maintains its computational advantages even as the complexity increases significantly.
Consider the eight-point CHY integrand
$I_{8}=\frac{1}{z_{12}^{3}z_{15}z_{26}z_{36}z_{38}^{3}z_{45}^{3}z_{47}z_{67}^{2}z_{78}},$ (49)
whose 4-regular graph representation is shown in Figure 39.
Figure 39: The 4-regular graph representation of the eight-point CHY integrand $I_{8}$. The edge multiplicities directly encode the denominator powers in equation (49). The triple edges at (1,2)(1,2)( 1 , 2 ), (3,8)(3,8)( 3 , 8 ), and (4,5)(4,5)( 4 , 5 ) immediately reveal three double poles, while the more complex edge structure suggests an additional higher-order pole that requires careful analysis.
Using the pole index formula, we identify all subsets with $\chi(A)>0$:
[…]
where $\chi(A)=1$ for each, indicating four constraints that must be addressed. The presence of the four-particle pole $s_{1245}$ makes this example particularly challenging, as it introduces correlations between the decomposition steps.
[…]
To reduce these higher-order poles and redundant items, we construct a six-point cross-ratio factor $\frac{z_{12}z_{38}z_{45}}{z_{18}z_{35}z_{42}}$ targeting the three two-particle double poles. Figure 40 shows its systematic decomposition.
Figure 40: Decomposition of the six-point cross-ratio factor for the eight-point integrand. The bipartite representation reveals the natural factorization into $\frac{z_{12}z_{38}}{z_{18}z_{32}}\times\frac{z_{38}z_{45}}{z_{35}z_{48}}$, providing the foundation for our two-stage decomposition strategy. This decomposition is specifically designed to address the three two-particle double poles while minimizing interference with the four-particle pole $s_{1245}$.
Round 1: Initial Decomposition
Based on the cross-ratio decomposition, we select two strategic identities:
$I_{8}[\{1,2\},1,8]$ $=-\frac{s_{23}z_{12}z_{38}}{s_{12}z_{18}z_{23}}-\frac{s_{24}z_{12}z_{48}}{s_{12}z_{18}z_{24}}-\frac{s_{25}z_{12}z_{58}}{s_{12}z_{18}z_{25}}-\frac{s_{26}z_{12}z_{68}}{s_{12}z_{18}z_{26}}-\frac{s_{27}z_{12}z_{78}}{s_{12}z_{18}z_{27}},$ (51)
$I_{8}[\{4,5\},5,3]$ $=-\frac{s_{41}z_{13}z_{45}}{s_{45}z_{14}z_{35}}-\frac{s_{42}z_{23}z_{45}}{s_{45}z_{24}z_{35}}-\frac{s_{46}z_{36}z_{45}}{s_{45}z_{35}z_{46}}-\frac{s_{47}z_{37}z_{45}}{s_{45}z_{35}z_{47}}-\frac{s_{48}z_{38}z_{45}}{s_{45}z_{35}z_{48}}.$ (52)
Multiplying $I_{8}$ by these identities produces $5^{2}=25$ terms, whose structure is shown in Figure 41.
Figure 41: Round 1 decomposition results: The initial application of identities (51) and (52) produces 25 terms organized in a $\times 5$ 5 grid. The first 9 terms (top-left, shown in gray) already contain only simple poles. The central 12 terms retain the double pole $s_{38}$, while the final 4 terms (bottom-right) contain both $s_{38}$ and $s_{1245}$ as double poles, requiring the most extensive further treatment.
The decomposition yields three categories:
  • 9 terms with only simple poles
  • 12 terms with double pole $s_{38}$
  • 4 terms with double poles $s_{38}$ and $s_{1245}$
Round 2: Eliminating the $s_{38}$ Double Pole
For the 12 terms containing double pole $s_{38}$, we apply the identity:
$I_{8}[\{3,8\},3,4]=-\frac{s_{81}z_{14}z_{38}}{s_{38}z_{18}z_{34}}-\frac{s_{82}z_{24}z_{38}}{s_{38}z_{28}z_{34}}-\frac{s_{85}z_{45}z_{38}}{s_{45}z_{34}z_{58}}-\frac{s_{86}z_{46}z_{38}}{s_{38}z_{34}z_{68}}-\frac{s_{87}z_{47}z_{38}}{s_{38}z_{34}z_{78}}.$ (53)
Figure 42 illustrates this decomposition for one representative term.
Figure 42: Round 2 decomposition of a term with double pole $s_{38}$. The application of identity (53) to the term $\frac{s_{24}s_{46}}{s_{12}s_{45}}\frac{z_{84}}{z_{21}^{2}z_{42}z_{51}z_{53}z_{54}^{2}z_{62}z_{64}z_{74}z_{76}^{2}z_{81}z_{83}^{3}z_{87}}$ produces 5 terms with only simple poles. The cross-ratio factors in the identity precisely cancel the $z_{83}^{3}$ contribution, reducing the pole order as required.
This process transforms the 12 terms with double pole $s_{38}$ into 12×5=6012$\times$ 5=6012 × 5 = 60 terms with only simple poles.
For the 4 terms containing both $s_{38}$ and $s_{1245}$ double poles, applying identity (53) yields 20 terms, of which 12 have only simple poles while 8 retain the double pole $s_{1245}$. Figure 43 shows a representative example.
Figure 43: Round 2 decomposition of a term with double poles $s_{38}$ and $s_{1245}$. After applying identity (53), the resulting 5 terms split into two categories: 3 terms (shown in gray) with only simple poles, and 2 terms (shown in white) that still contain the double pole $s_{1245}$, necessitating a third round of decomposition.
Round 3: Final Decomposition of $s_{1245}$
The remaining 8 terms with double pole $s_{1245}$ require the application of a nine-term identity:
$I_{8}[\{1,2,4,5\},1,3]=$ $-\frac{s_{26}z_{12}z_{36}}{s_{1245}z_{13}z_{26}}-\frac{s_{46}z_{14}z_{36}}{s_{1245}z_{13}z_{46}}-\frac{s_{56}z_{15}z_{36}}{s_{1245}z_{13}z_{56}}-\frac{s_{27}z_{12}z_{37}}{s_{1245}z_{13}z_{27}}$
$-\frac{s_{28}z_{12}z_{38}}{s_{1245}z_{13}z_{28}}-\frac{s_{47}z_{14}z_{37}}{s_{1245}z_{13}z_{47}}-\frac{s_{48}z_{14}z_{38}}{s_{1245}z_{13}z_{48}}-\frac{s_{57}z_{15}z_{37}}{s_{1245}z_{13}z_{57}}-\frac{s_{58}z_{15}z_{38}}{s_{1245}z_{13}z_{58}}.$ (54)
Figure 44 demonstrates the final decomposition step.
Figure 44: Round 3 final decomposition. The nine-term identity (54) applied to a term with double pole $s_{1245}$ produces 9 CHY integrands with only simple poles. This completes the systematic reduction of all higher-order poles in the original eight-point integrand. The complexity of this final step (9 terms) reflects the four-particle nature of the $s_{1245}$ pole.
Through this three-round process, the eight-point CHY integrand $I_{8}$ with four double poles is systematically decomposed into 9+60+12+72=1539+60+12+72=1539 + 60 + 12 + 72 = 153 terms, each containing only simple poles. The improvement over traditional methods is summarized in the following table:
Method Round 1 Round 2 Round 3 Round 4 Total
Systematic algorithm Cardona:2016gon 5 [5] 25 [20] 105 [20] 265 [0] 265
Our method 25 [16] 89 [8] 153 [0] 153
Numbers in brackets [·] indicate terms with remaining higher-order poles at each stage.